### v2

# Large $N$ Wess-Zumino model at finite temperature and large chemical potential in $3d$

Abstract
We consider the supersymmetric Wess-Zumino model at large $N$ in $(2+1)$ dimension. We introduce a chemical potential($\mu$) at finite temperature($T$). The non-trivial fixed point of this model is described by a pair of coupled gap equations. This fixed point behaves as a thermal CFT for all values of the coupling. We find that at large chemical potential these coupled equations simplify and solutions become analytically tractable. We solve them analytically for all values of the coupling at this limit. The solutions admit a systematic series expansion in $\frac{T}{\mu}$. Thus, using the solutions of the gap equation at large chemical potential we can evaluate the analytic form of the partition function, stress tensor and spin-1 current as a perturbative expansion in orders of $\frac{T}{\mu}$. Applying the OPE inversion formula on the scalar and fermion two point functions of the theory, we compute higher spin currents at large $\mu.$
1 Introduction
The Wess-Zumino model was introduced as one of the simplest examples of supersymmetric quantum field theory in the literature Wess:1974tw , see for review Moshe:2003xn . The model consists of equal number of bosonic and fermionic degrees of freedom and it allows both intra-species as well as inter-species interactions. The two kinds of couplings are governed by a single coupling constant. Thermodynamics of this model in $(2+1)$ dimensions was studied in DeWolfe:2019etx at large $N$ and at zero chemical potential. The theory behaves as a thermal CFT for any value of the coupling constant and the conformal fixed point is characterised by the bosonic and fermion thermal masses satisfying a pair of coupled gap equations. And the thermodynamic quantities e.g. pressure, energy density had been evaluated by solving the coupled gap equations for the thermal masses numerically.
We introduce a chemical potential $\mu$ in this model starting from the supersymmetric Wess-Zumino model without chemical potential in the fundamental of $O(2N)$. Rewriting this in terms of complex scalars and fermions we get the $U(N)$ invariant action with complex scalars and Dirac fermions in dimension $d=3$ with the Wess-Zumino interaction. Now the action can be linearised using the Hubbard-Stratanovich trick introducing auxiliary fields at large $N$. Only the zero mode of the auxiliary field contributes at the leading order. Finally we evaluate the partition function at the saddle point of the path integral in the zero mode of the auxiliary fields. Two saddle point conditions give the pair of gap equations involving the chemical potential $\mu$. Now the modifications in the gap equations caused by the presence of chemical potential $\mu$ is simple to understand. The pair of coupled gap equation describing the conformal fixed point of the theory at any value of the coupling constant $\lambda$ is given by
$\frac{m_{f}}{4\lambda}$ $=-\frac{m_{b}}{2\pi}-\frac{T}{2\pi}[\log(1-e^{\frac{\mu-m_{b}}{T}})+\log(1-e^{-\frac{m_{b}+\mu}{T}})],$
$\frac{m_{b}^{2}}{8\lambda}$ $=\frac{3m_{f}^{2}}{8\lambda}+\frac{m_{f}^{2}}{2\pi}+\frac{m_{f}T}{2\pi}[\log(1+e^{\frac{\mu-m_{f}}{T}})+\log(1+e^{-\frac{m_{f}+\mu}{T}})],$ (1)
where $m_{b}$ and $m_{f}$ are thermal masses of bosons and fermions respectively, $T$ is the temperature.
We observe that at large chemical potential $\frac{\mu}{T}\to\infty$, the gap equations (1) simplifies and can be solved at this limit. The solutions admit systematic series expansion in $\frac{T}{\mu}$. At the leading order in large $\mu$ in units of temperature($T$), the solutions of the gap equation for the thermal mass of scalars and fermions are given as
$m_{b}=\mu,\qquad m_{f}=\frac{(\sqrt{4\lambda^{2}+3\pi^{2}}-2\lambda)\mu}{3\pi}.$ (2)
Study of thermal one point functions of CFT operators has emerged as an important question in recent developments in the context of holography as well as finite temperature CFTs. In this paragraph, we provide a summary of the recent advancements in this field. It has been shown that holographic CFT thermal one point functions can probe the interior geometry of the $AdS$ black hole in the bulk Grinberg:2020fdj ; Berenstein:2022nlj ; David:2022nfn ; Singhi:2024sdr . Evaluating thermal one point function from field theory turns out to be a challenging task. In this context a class of large $N$ vector models serve as a suitable examples of CFT’s at finite temperature where thermal one point functions of higher spin currents are computed based on the simplification at large $N$. Methods to evaluate thermal one point functions of arbitrary spin($l$) currents has been developed for $O(N)$ model at large $N$ examining the OPE of the two point functions at the non-trivial fixed point Sachdev:1993pr ; Chubukov:1993aau of the theory, the method is called OPE inversion formula Iliesiu:2018fao . At zero temperature the model in $d=3$ is dual to the higher spin gravity theory in the bulk $AdS_{4}$ geometry Klebanov:2002ja . The OPE inversion formula has been generalised for large $N$ Gross-Neveu model for fermions Petkou:2018ynm ; David:2023uya . Thermal one point functions of higher spin currents in presence of chemical potential in the theory are computed by applying the OPE inversion formula on the two point functions of twisted fields in Karydas:2023ufs ; David:2024naf . For these large $N$ vector models the leading $\frac{1}{N}$ corrections to the free energy as well as the thermal one point functions are discussed in Diatlyk:2023msc . Bootstrap techniques are also developed to evaluate thermal OPE coefficients of the $O(N)$ model containing the thermal one point functions of higher spin currents at finite $N$ Barrat:2024fwq . We have computed thermal one point functions of CFT higher spin currents for $O(N)$ model on $S^{1}\times S^{2}$ in David:2024pir as a perturbative expansion at large radius of the sphere $S^{2}$. These one point functions of higher spin currents show interesting behaviour both at large spin($l$) and large space time dimension($d$) David:2023uya ; David:2024naf . For any general CFT on $S^{1}\times S^{2}$ the one point functions are examined in Buric:2024kxo .
In this paper, we will compute thermal one point functions of the operators constructed out of scalar and fermion bilinears for the large $N$ Wess-Zuminio model with chemical potential $\mu$ in $3d$ applying the OPE inversion formula. Finally, incorporating the solutions of the gap equations at large $\mu$ given in (2), in these thermal one point functions we obtain their large $\mu$ behaviour. We study thermal one point functions of the classes of operators in this Wess-Zumino model at the large $N$ conformal fixed point given as follows.
$\langle{\cal O}[n,l]\rangle$ $\equiv\langle\phi^{*}D_{\mu_{1}}\cdots D_{\mu_{l}}D^{2n}\phi\rangle\equiv b_{\cal O}[n,l]T^{\Delta_{\cal O}},$
$\langle\tilde{{{\cal O}}}_{0}[n,l]\rangle$ $\equiv\langle\psi_{\alpha}^{\dagger}D_{\mu_{1}}\cdots D_{\mu_{l}}D^{2n}\psi_{\alpha}\rangle\equiv b_{{\cal\tilde{O}}_{0}}[n,l]T^{\Delta_{{\cal\tilde{O}}_{0}}},$
$\langle\tilde{{{\cal O}}}_{+}[n,l]\rangle$ $\equiv\langle\psi_{\alpha}^{\dagger}\gamma^{\alpha\beta}_{\mu_{1}}D_{\mu_{2}}\cdots D_{\mu_{l}}D^{2n}\psi_{\beta}\rangle\equiv b_{{\cal O}_{+}}[n,l]T^{\Delta_{{\cal\tilde{O}}_{-}}},$
$\langle\tilde{{{\cal O}}}_{-}[n,l]\rangle$ $\equiv\langle\psi_{\alpha}^{\dagger}\gamma^{\alpha\beta}_{\mu}D^{\mu}D_{\mu_{1}}\cdots D_{\mu_{l}}D^{2n}\psi_{\beta}\rangle\equiv b_{{\cal O}_{-}}[n,l]T^{\Delta_{{\cal\tilde{O}}_{+}}}.$ (3)
where $D_{\nu}=\partial_{\nu}+\mu\delta_{\nu,0}$ with $\nu=0,1,2$ and $\gamma_{\mu}^{\alpha\beta}$ denotes the Dirac gamma matrices. These bilinear operators in the above four equations are understood to be traceless symmetric parts of the objects written above. $\Delta_{{\cal O}},\Delta_{{\cal O}_{0}},\Delta_{{\cal O}_{+}},\Delta_{{\cal O}_{-}}$ are the scaling dimensions of the corresponding operators.
Notation: Throughout this paper the notation $\tilde{\cal O}$ stands for the operators constructed of the fermion bilinears while $\cal O$ denotes scalar bilinears.
The thermal expectation value of these operators given in (1) are evaluated by studying the OPEs of the thermal two point functions of complex scalars and fermions simultaneously with the use of OPE inversion formula. We allow the fermion bilinear to occur in the OPE of the scalar two point correlator and vice versa. The reason is that the interactions between scalar and fermion fields in the action may cause non-trivial three point function coefficient between scalar, scalar, fermionic bilinear operator and also fermion, fermion, scalar bilinear operators. Starting from such a general structure of these OPEs we see that such mixing of operators are not present at the leading order in large $N$. The three-point function between fermion, fermion, and scalar bilinear comes out to be zero from the OPE inversion formula directly. Although for scalar, scalar, fermion bilinear it does not come out directly from the OPE inversion formula, we argue that this three point function is zero based on the fact that it is known to vanish at the Gaussian fixed point as well as at the non-trivial fixed point obtained at strong coupling. We rigorously show that it indeed vanishes for $l=1$ and $2$. Once we obtain the thermal one point functions in terms of the thermal masses of bosons and fermions we take the large $\mu$ limit of the results with the use of the large $\mu$ solutions of the gap equations (2). At large $\mu$ we get the thermal one point function of the scalar bilinear ${\cal O}[0,l]$ as a systematic series expansion as given below
$\lim_{\mu\rightarrow\infty}a^{b}_{\cal O}[0,l]=\frac{\mu^{l+1}\Gamma(l+1)}{2^{2-l}\pi\Gamma(2l+1)}\Bigg{[}1+\frac{\sqrt{4\lambda^{2}+3\pi^{2}}-2\lambda}{6\lambda}+$ $\sum_{n=1}^{l}\frac{(l-n+1)_{2n}\zeta(n+1)}{2^{n}n!}\frac{T^{n+1}}{|\mu|^{n+1}}\Bigg{]}.$ (4)
The corrections appear as $O(e^{-\frac{\mu}{T}})$. The above expression also contains overall constant factors along with $\langle{\cal O}[0,l]\rangle$ due to the OPE coefficient which does not depend on the dynamics at finite temperature. And at large $\mu$ the fermion bilinear operator ${\cal O}_{+}[0,l]$ has thermal expectation given by the following expansion
$\lim_{\mu\to\infty}a_{{\cal\tilde{O}}_{+}}[0,l]=-\frac{l(c\mu)^{l+1}}{\pi 2^{l+1}}\sum_{k=0}^{l+1}\frac{(2\pi i)^{k}B_{k}\left(\frac{1}{2}\right)}{\left(\frac{1}{2}\right)_{l}(2c)^{k}}\Big{(}\frac{T}{\mu}\Big{)}^{k}\frac{\,{}_{2}\tilde{F}_{1}\left(k-l-1,k+l;k;\frac{c-1}{2c}\right)}{k!(k+l)_{2-2k}},$
${\rm where}\qquad c=\frac{(\sqrt{4\lambda^{2}+3\pi^{2}}-2\lambda)}{3\pi}.$ (5)
This expression also has an overall constant factor due to the OPE coefficients independent of finite temperature dynamics. Note that this expression is always real as $B_{k}(\frac{1}{2})$ vanishes for $k$ being odd. Again, the corrections occur in order $O(e^{-\frac{\mu}{T}})$.
The organisation of the paper is as follows. In the section 2 we discuss general structures of the OPEs for scalar and fermion two point functions for theories with scalar-fermion interactions. The section 3 contains the description of the Wess-Zumino model at finite chemical potential and its analytic solutions at large $N$ and large chemical potential. Section 4 has discussions of the OPE inversion formula and the results for the thermal one point functions. Section 5 has the large $\mu$ asymptotic formula for the thermal one point functions. The agreement of the spin-1 and spin-2 current with the partition function calculation is discussed in section 6. The appendix A has the details of regularisation used to evaluate the partition function for free scalar and free fermions.
2 OPEs in theories with both bosons and fermions
In this section we introduce the operator product expansion of the two point correlation functions of scalars and fermions. It is essential to examine both of these while studying the models constituted of both bosonic and fermion fields. Important and non-trivial modifications arise due to the inter-species interaction between scalar and fermionic fields, which allows the possibility of bosonic two point function decomposing into fermionic bilinear operators in the OPE and as well as the two point correlator of fermion fields admitting the presence of the bosonic bilinear operators in its OPE. Identifying all sorts operators present in the OPEs of both scalar and fermion two point functions, we organise these OPEs in terms of Gegenbauer polynomials. This sets the stage for the computation of the thermal one point functions for all theses operators constructed from scalar and fermion bilinears by the application of OPE inversion formula at finite temperature.
2.1 Scalar two point function
We consider theories with both bosons and fermions and allow interaction between them. The general structure of OPEs for the two point function of scalars $\phi$ may contain fermionic bilinears owing to such interactions as three point function of boson, boson and fermionic bilinears can be non-vanishing in general.
$\phi^{*}(x)\phi(0)=\sum_{\mathcal{O}}C_{\phi\phi\mathcal{O}}|x|^{-2\Delta_{\phi}+\Delta_{\mathcal{O}}-l}x_{\mu_{1}}\cdots x_{\mu_{l}}\mathcal{O}^{\mu_{1}\cdots\mu_{l}}$
$+\sum_{\mathcal{\tilde{O}}}C_{\phi\phi\gamma^{\alpha\beta}{\cal\tilde{O}}_{\alpha\beta}}|x|^{-2\Delta_{\phi}+\Delta_{\tilde{\mathcal{O}}}-l-1}x_{\mu_{1}}\cdots x_{\mu_{l}}x_{\mu}\gamma^{\alpha\beta\,\mu}\tilde{\mathcal{O}}^{\mu_{1}\cdots\mu_{l}}_{\alpha\beta}$
$+\sum_{\mathcal{\tilde{O}}}C_{\phi\phi{\cal\tilde{O}}_{\alpha\alpha}}|x|^{-2\Delta_{\phi}+\Delta_{\tilde{\mathcal{O}}}-l}x_{\mu_{1}}\cdots x_{\mu_{l}}\tilde{\mathcal{O}}^{\mu_{1}\cdots\mu_{l}}_{\alpha\alpha},$ (6)
$\mathcal{O}$ and $O_{\alpha\beta}$ refer to the bosonic and fermionic bilinear operators respectively. $\alpha,\beta$ stand for the spinor indices due to the fermionic fields and $\mu_{1},\cdots,\mu_{l}$ are space-time indices, repeated indices are understood to be summed over. $C_{\phi\phi{\cal O}},C_{\phi\phi\gamma^{\alpha\beta}{\cal O}_{\alpha\beta}}$ and $C_{\phi\phi{\cal O}_{\alpha\alpha}}$ are the relevant OPE coefficients which are the ratio of the structure constant to the normalisation of the operators appearing in the OPE. The bosonic and fermionic bilinear operators arising in the above OPE have the schematic forms as given below,
$\mathcal{O}^{\mu_{1}\cdots\mu_{l}}\equiv\phi^{*}\partial_{\mu_{1}}\cdots\partial_{\mu_{l}}(\partial^{2})^{n}\phi,$
$\tilde{\mathcal{O}}_{\alpha\beta}^{\mu_{1}\cdots\mu_{l}}\equiv\psi_{\alpha}^{\dagger}\partial_{\mu_{1}}\cdots\partial_{\mu_{l}}(\partial^{2})^{n}\psi_{\beta}.$ (7)
Now the spatial rotational and spatial translational invariance of the thermal vacuum fix the structure of the thermal one point functions of bosonic bilinears except a numerical constant to be determined. Thus thermal one point functions of the class of bosonic operators denoted by ${\cal O}$ can be represented as traceless symmetric tensors as given below,
$\langle\mathcal{O}_{\mu_{1}\cdots\mu_{l}}\rangle=b_{\mathcal{O}}T^{\Delta}(e_{\mu_{1}}e_{\mu_{2}}\cdots e_{\mu_{l}}-\text{traces}),$ (8)
$e_{\mu}$’s are unit vectors along imaginary time direction $\tau$. $b_{\mathcal{O}}$ is the numerical constant specific to a theory. And using the following mathematical identity,
$|x|^{-l}(x^{\mu_{1}}\cdots x^{\mu_{l}})(e_{\mu_{1}}\cdots e_{\mu_{l}}-\text{traces})=\frac{l!}{2^{l}(\nu)_{l}}C_{l}^{(\nu)}(\eta),$ (9)
where $C_{l}^{\nu}(\eta)$ is Gegenbauer polynomial with
$\eta=\frac{\tau}{|x|},\qquad{\rm and}\qquad\nu=\frac{d-2}{2}.$ (10)
The 1st term from (2.1) is expressed in terms of Gegenbauer polynomials,
$\sum_{\mathcal{O}}C_{\phi\phi\mathcal{O}}|x|^{-2\Delta_{\phi}+\Delta_{\mathcal{O}}-l}x_{\mu_{1}}\cdots x_{\mu_{l}}\mathcal{O}^{\mu_{1}\cdots\mu_{l}}=\sum_{\mathcal{O}}a_{\cal O}^{b}|x|^{-2\Delta_{\phi}+\Delta_{\mathcal{O}}}C_{l}^{(\nu)}(\eta),$ (11)
where,
$a_{\cal O}^{b}=C_{\phi\phi\mathcal{O}}\frac{b_{\mathcal{O}}l!}{2^{l}(\nu)_{l}}T^{\Delta_{\cal O}}.$ (12)
Here the notation $b$ in the superscript means that it is derived from the OPE for the bosonic 2-point function. This is used to specify the kind of structure constant multiplied with the one point functions.
Now we focus on the 2nd term from the OPE (2.1), and the fermionic bilinear operator appearing in this term is decomposed into a symmetric part, trace part and an antisymmetric part, as was done in David:2023uya .
$\gamma^{\nu\;\alpha\beta}{\cal\tilde{O}}^{\mu_{1}\cdots\mu_{l}}_{\beta\alpha}$ $=$ $\left[\frac{1}{l+1}\left(\gamma^{\nu\;\alpha\beta}{\cal\tilde{O}}^{\mu_{1}\cdots\mu_{l}}_{\beta\alpha}+\gamma^{\mu_{1}\;\alpha\beta}{\cal\tilde{O}}^{\mu_{2}\cdots\mu_{l}\nu}_{\beta\alpha}+{\rm cyclic}\right)-{\rm Traces}\right]+{\rm Traces}$ (13)
$+\frac{1}{l+1}\left(\gamma^{\nu\;\alpha\beta}{\cal\tilde{O}}^{\mu_{1}\cdots\mu_{l}}_{\beta\alpha}-\gamma^{\mu_{1}\;\alpha\beta}{\cal\tilde{O}}^{\nu\mu_{2}\cdots\mu_{l}}_{\beta\alpha}\right)+\cdots(l-1){\rm terms},$
The ‘Traces’ stands for
${\rm Traces}=\frac{2}{(l+1)(d+l-1)}\left(\delta^{\mu\mu_{1}}\gamma^{\rho\;\alpha\beta}{\cal O}^{\rho_{\mu_{2}}\cdots\mu_{l}}_{\beta\alpha}+\delta^{\mu_{1}\mu_{2}}\gamma^{\rho\;\alpha\beta}{\cal O}^{\rho_{\mu_{3}}\cdots\mu_{l}\mu_{1}}_{\beta\alpha}+{\rm cyclic}\right).$ (14)
Finally, using the mathematical identity (9) after substituting this tensor structure (13), the 2nd term from the OPE (2.1) is simplified to,
$\sum_{\mathcal{\tilde{O}}}C_{\phi\phi\gamma^{\alpha\beta}{\cal O}_{\alpha\beta}}|x|^{-2\Delta_{\phi}+\Delta_{\tilde{\mathcal{O}}}-l-1}x_{\mu_{1}}\cdots x_{\mu_{l}}x_{\mu}\gamma^{\alpha\beta\,\mu}{\mathcal{O}}^{\mu_{1}\cdots\mu_{l}}_{\alpha\beta}$
$=\sum_{\tilde{\cal O}}|x|^{\Delta_{{\cal\tilde{O}}}-2\Delta_{\phi}}\left(a_{{\cal\tilde{O}}_{+}}C^{(\nu)}_{l+1}(\eta)+\frac{2l}{(l+1)(d+l-1)}a_{{\cal\tilde{O}}_{-}}C^{(\nu)}_{l-1}(\eta)\right),$ (15)
where,
$a_{{\cal\tilde{O}}_{+}}^{b}=b_{{\cal O}_{+}}T^{\Delta_{{\cal O}_{+}}}\frac{(l+1)!}{2^{l+1}(\nu)_{l+1}}C_{\phi\phi\gamma^{\alpha\beta}{\cal O}_{\alpha\beta}},$ (16)
$a_{{\cal\tilde{O}}_{-}}^{b}=b_{{\cal O}_{-}}T^{\Delta_{{\cal O}_{-}}}\frac{(l-1)!}{2^{l-1}(\nu)_{l-1}}C_{\phi\phi\gamma^{\alpha\beta}{\cal O}_{\alpha\beta}}.$
Now in the 3rd term in the OPE (2.1) the fermionic operator of the class $\tilde{\cal O}_{0}$ again can be written as traceless symmetric tensor as follows,
$\langle\mathcal{O}_{\alpha\alpha}^{\mu_{1}\cdots\mu_{l}}\rangle_{\beta}=b_{{\cal O}_{0}}T^{\Delta}(e_{\mu_{1}}e_{\mu_{2}}\cdots e_{\mu_{l}}-\text{traces}),$ (17)
And also the third term from (2.1) is written as,
$\sum_{\mathcal{\tilde{O}}}C_{\phi\phi{\cal\tilde{O}}_{\alpha\alpha}}|x|^{-2\Delta_{\phi}+\Delta_{\tilde{\mathcal{O}}}-l}x_{\mu_{1}}\cdots x_{\mu_{l}}{\mathcal{\tilde{O}}}^{\mu_{1}\cdots\mu_{l}}_{\alpha\alpha}=\sum_{\mathcal{\tilde{O}}}a_{{\cal\tilde{O}}_{0}}^{b}|x|^{-2\Delta_{\phi}+\Delta_{\mathcal{\tilde{O}}}}C_{l}^{(\nu)}(\eta)$
where
$a_{{\cal\tilde{O}}_{0}}^{b}=C_{\phi\phi{\cal\tilde{O}}_{\alpha\alpha}}\frac{b_{\mathcal{\tilde{O}}_{0}}l!}{2^{l}(\nu)_{l}}.$ (18)
Thus combining all the three terms in the OPE (2.1) as obtained in (11), (2.1) and (2.1) respectively, and if the anomalous dimensions of the operators present in the OPE are vanishing at the leading order in large $N$ [^1] Our analysis in the section 4 will establish the fact that for large $N$ Wess-Zumino model in $3d$ operator dimensions do not get corrected at the leading order in large $N$., we can use
$\Delta_{\cal O}=2\Delta_{\phi}+2n+l\qquad{\rm and}\qquad\Delta_{\cal\tilde{O}}=2\Delta_{\psi}+2n+l,$
${\rm where}\qquad\Delta_{\phi}=\frac{d-2}{2}\qquad{\rm and}\qquad\Delta_{\psi}=\frac{d-1}{2}.$ (19)
And present the OPE (2.1) in terms of the Gegenbauer polynomials as
$\langle\phi(x)\phi(0)\rangle=$ $\sum_{n,l=0}^{\infty}a_{\mathcal{O}}^{b}|x|^{2n+l}C_{l}^{\nu}(\eta)+\sum_{l,n=0}^{\infty}|x|^{2n+l+1}a_{{\cal\tilde{O}}_{+}}^{b}C^{(\nu)}_{l+1}(\eta)$ (20)
$+\sum_{l=1,n=0}^{\infty}|x|^{2n+l+1}\frac{2l}{(l+1)(l+d-1)}a_{{\cal\tilde{O}}_{-}}^{b}C^{(\nu)}_{l-1}(\eta)+\sum_{n,l=0}^{\infty}a_{\mathcal{\tilde{O}}_{0}}^{b}|x|^{2n+l+1}C_{l}^{\nu}(\eta)$
Reorganising the sums in the above equation we obtain the OPE as,
$\langle\phi(x)\phi(0)\rangle=\sum_{n=1}^{\infty}a_{\mathcal{O}}^{b}[n,0]|x|^{2n}C_{0}^{(\nu)}(\eta)+\sum_{l=1}^{\infty}(a^{b}_{\mathcal{O}}[0,l]+a^{b}_{{\cal\tilde{O}}_{+}}[0,l])|x|^{l}C_{l}^{(\nu)}(\eta)$
$+\sum_{l=0,n=1}^{\infty}|x|^{2n+l}\Big{(}a^{b}_{\mathcal{O}}[n,l]+a^{b}_{{\cal\tilde{O}}_{+}}[n,l]+\frac{2(l+1)}{(l+2)(d+l)}a^{b}_{{\cal\tilde{O}}_{-}}[n-1,l]\Big{)}C^{(\nu)}_{l}(\eta)$
$+\sum_{n,l=0}^{\infty}a^{b}_{\mathcal{\tilde{O}}_{0}}[n,l]|x|^{2n+l+1}C_{l}^{(\nu)}(\eta).$ (21)
The above structure will be useful in the analysis of the OPE inversion formula in the Section 4.
2.2 Fermion two point functions
Similar to the bosonic case, due to the fermion-boson interaction the OPE of the fermionic two point function allows presence of bosonic bilinear operators as the three point function of fermion, fermion and bosonic bilinear can be non-vanishing in general just as we had similar considerations for the bosonic OPE. Thus combining all sorts of operators allowed, the general structure for the OPE of the fermionic two point function $\psi_{\alpha}(x)$ can be written as,
$g_{\alpha\beta}(x)=\langle\psi^{\dagger}_{\alpha}(0)\psi_{\beta}(x)\rangle=\sum_{\mathcal{\tilde{O}}}C_{\psi\psi\mathcal{\tilde{O}}}|x|^{-2\Delta_{\psi}+\Delta_{\mathcal{\tilde{O}}}-l}x_{\mu_{1}}\cdots x_{\mu_{l}}\langle\mathcal{\tilde{O}}_{\mu_{1}\cdots\mu_{l}}^{\alpha\beta}\rangle$
$+\sum_{\mathcal{O}}C_{\psi\psi\mathcal{O}}|x|^{-2\Delta_{\psi}+\Delta_{\mathcal{O}}-l-1}x_{\mu_{1}}\cdots x_{\mu_{l}}x^{\mu}\gamma^{\alpha\beta}_{\mu}\langle\mathcal{O}_{\mu_{1}\cdots\mu_{l}}\rangle$
$+\sum_{\mathcal{O}}C_{\psi\psi\mathcal{O}}|x|^{-2\Delta_{\psi}+\Delta_{\mathcal{O}}-l}x_{\mu_{1}}\cdots x_{\mu_{l}}\delta^{\alpha\beta}\langle\mathcal{O}_{\mu_{1}\cdots\mu_{l}}\rangle.$ (22)
The 1st term in the above operator product expansion of fermion two point correlator has the bilinear operator of fermion with spinor indices $\alpha,\beta$ while the 2nd and 3rd terms are decompositions of the fermion two point function into bosonic bilinears through two different projection channels in the OPE. And $C_{\psi\psi\cal\tilde{O}},C_{\psi\psi\cal O}$ and $C_{\psi\psi\cal O}$ refer to the relevant OPE coefficients.
Now, we define three different correlators of fermion field $g_{1}(x),g_{2}(x),g_{3}(x)$ obtained by projecting the $g_{\alpha\beta}(x)$ in three different channels as shown below,
$g_{1}(x)$ $=\langle\psi_{\alpha}^{\dagger}(x)\psi_{\alpha}(0)\rangle_{\beta},$
$g_{2}(x)$ $=\frac{1}{|x|}\langle\gamma^{\alpha\beta}_{\mu}x^{\mu}\psi_{\alpha}^{\dagger}(x)\psi_{\beta}(0)\rangle_{\beta},$
$g_{3}(x)$ $=\langle\gamma_{\mu}^{\alpha\beta}\partial^{\mu}\psi_{\alpha}^{\dagger}(x)\psi_{\alpha}(0)\rangle_{\beta}.$ (23)
By studying these three OPE channels simultaneously, using OPE inversion formula the thermal one point functions of different classes of both bosonic and fermionic bilinears can be computed separately.
The correlator $g_{1}(x)$
Form (2.2) and (2.2) the OPE for the correlator $g_{1}(x)$ is given by,
$g_{1}(x)=\langle\psi^{\dagger}_{\alpha}(x)\psi_{\alpha}(0)\rangle$ $=\sum_{\mathcal{\tilde{O}}}C_{\psi\psi\mathcal{\tilde{O}}}|x|^{-2\Delta_{\psi}+\Delta_{\mathcal{\tilde{O}}}-l}x^{\mu_{1}}\cdots x^{\mu_{l}}\langle\mathcal{\tilde{O}}_{\mu_{1}\cdots\mu_{l}}^{\alpha\alpha}\rangle$
$\qquad+2\sum_{\mathcal{O}}C_{\psi\psi\mathcal{O}}|x|^{-2\Delta_{\psi}+\Delta_{\mathcal{O}}-l}x^{\mu_{1}}\cdots x^{\mu_{l}}\langle\mathcal{O}_{\mu_{1}\cdots\mu_{l}}\rangle$ (24)
$=\sum_{\mathcal{O}}a^{f}_{\mathcal{\tilde{O}}_{0}}|x|^{-2\Delta_{\psi}+\Delta_{\mathcal{\tilde{O}}}}C_{l}^{(\nu)}(\eta)+2\sum_{\mathcal{O}}a^{f}_{\mathcal{O}}|x|^{-2\Delta_{\psi}+\Delta_{\mathcal{O}}}C_{l}^{(\nu)}(\eta),$ (25)
Where
$a^{f}_{{\cal\tilde{O}}_{0}}=C_{\psi\psi\mathcal{\tilde{O}}}\frac{b_{\mathcal{\tilde{O}}_{0}}l!}{2^{l}(\nu)_{l}}T^{\Delta_{\cal O}}\qquad{\rm and}\qquad a^{f}_{\cal O}=C_{\psi\psi{\cal O}}\frac{b_{\mathcal{O}}l!}{2^{l}(\nu)_{l}}T^{\Delta_{\cal O}}.$ (26)
Certainly the notation $f$ is used to identify the thermal one point functions of the operators derived from the OPE of fermion 2-point correlators. Similar to the case of scalars we use (2.1) to obtain
$g_{1}(x)=\sum_{n,l=0}^{\infty}a_{\mathcal{\tilde{O}}_{0}}^{f}[n,l]|x|^{2n+l}C_{l}^{(\nu)}(\eta)+2\sum_{n,l=0}a^{f}_{\mathcal{O}}[n,l]|x|^{2n+l-1}C_{l}^{(\nu)}(\eta).$ (27)
The correlator $g_{2}(x)$
Projecting the OPE (2.2) in the vector channel using (2.2) we can compute $g_{2}(x)$ as the following
$g_{2}(x)=\langle g_{\alpha\beta}\frac{\gamma^{\alpha\beta}_{\nu}x^{\nu}}{|x|}\rangle=\sum_{\mathcal{\tilde{O}}}C_{\psi^{\dagger}\psi\mathcal{\tilde{O}}}|x|^{\Delta_{\mathcal{\tilde{O}}}-2\Delta_{\psi}-l-1}x^{\mu_{1}}\cdots x^{\mu_{l}}x^{\nu}\langle\gamma^{\alpha\beta}_{\nu}\mathcal{\tilde{O}}_{\alpha\beta\mu_{1}\cdots\mu_{l}}\rangle$
$+\sum_{\mathcal{O}}C_{\psi^{\dagger}\psi\mathcal{O}}|x|^{\Delta_{\mathcal{O}}-2\Delta_{\psi}-l}\,x^{\mu_{1}}\cdots x^{\mu_{l}}\langle\mathcal{O}_{\mu_{1}\cdots\mu_{l}}\rangle.$ (28)
The 1st term from the above equation can be organised as the following, using the similar relation as (2.1),
$\sum_{\mathcal{\tilde{O}}}C_{\psi^{\dagger}\psi\mathcal{\tilde{O}}}|x|^{\Delta_{\mathcal{\tilde{O}}}-2\Delta_{\psi}-l-1}x^{\mu_{1}}\cdots x^{\mu_{l}}x^{\nu}\langle\gamma^{\alpha\beta}_{\nu}\mathcal{\tilde{O}}_{\alpha\beta\mu_{1}\cdots\mu_{l}}\rangle$
$=\sum_{{\mathcal{\tilde{O}}}}|x|^{\Delta_{\mathcal{\tilde{O}}}-2\Delta_{\psi}}\left(a^{f}_{\mathcal{\tilde{O}}_{+}}C^{(\nu)}_{l+1}(\eta)+\frac{2l}{(l+1)(d+l-1)}a^{f}_{\mathcal{\tilde{O}}_{-}}C^{(\nu)}_{l-1}(\eta)\right),$ (29)
where
$a_{{\cal\tilde{O}}_{+}}^{f}=b_{{\cal O}_{+}}T^{\Delta_{{\cal O}_{+}}}\frac{(l+1)!}{2^{l+1}(\nu)_{l+1}}C_{\psi\psi{\cal\tilde{O}}},$ (30)
$a_{{\cal\tilde{O}}_{-}}^{f}=b_{{\cal O}_{-}}T^{\Delta_{{\cal O}_{-}}}\frac{(l-1)!}{2^{l-1}(\nu)_{l-1}}C_{\psi\psi{\cal\tilde{O}}}.$
Thus, substituting this in (2) we have the fermionic correlator $g_{2}$ given by the following
$g_{2}(x)=\sum_{{\mathcal{\tilde{O}}}}|x|^{\Delta_{\mathcal{\tilde{O}}}-2\Delta_{\psi}}\left(a^{f}_{\mathcal{\tilde{O}}_{+}}C^{(\nu)}_{l+1}(\eta)+\frac{2l}{(l+1)(d+l-1)}a^{f}_{\mathcal{\tilde{O}}_{-}}C^{(\nu)}_{l-1}(\eta)\right)$
$+\sum_{\mathcal{O}}|x|^{\Delta_{\mathcal{O}}-2\Delta_{\psi}}a^{f}_{\mathcal{O}}C_{l}^{(\nu)}(\eta).$ (31)
The above OPE is reorganised as the following with the use of (2.1)
$g_{2}(x)$ $=\frac{a^{f}_{\mathcal{O}}[0,0]}{|x|}C_{0}^{(\nu)}(\eta)+\sum_{l=1}^{\infty}(a^{f}_{\mathcal{\tilde{O}}_{+}}[0,l]+a^{f}_{\mathcal{O}}[0,l])|x|^{l-1}C_{l}^{(\nu)}(\eta)$
$+\sum_{n=1,l=1}^{\infty}\Big{(}a^{f}_{\mathcal{\tilde{O}}_{+}}[n,l]+a^{f}_{\mathcal{O}}[n,l]+\frac{2(l+1)}{(l+2)(d+l)}a^{f}_{\mathcal{\tilde{O}}_{-}}[n-1,l+1]\Big{)}|x|^{2n+l-1}C_{l}^{(\nu)}(\eta)$ (32)
The Correlator $g_{3}(x)$
Using the similar approach, from (2.2) and (2.2), it is easy to evaluate the correlator $g_{3}(x)$ to be given by the following expression
$g_{3}(x)=$ $\sum_{\mathcal{\mathcal{O}}}|x|^{\Delta_{\mathcal{O}}-2\Delta_{\psi}-1}(\Delta_{\mathcal{O}}-2\Delta_{\psi}+2)a_{\mathcal{O}}[n,l]C_{l}^{(\nu)}(\eta)$
$+\sum_{\mathcal{\mathcal{O}}}|x|^{\Delta_{\mathcal{O}}-2\Delta_{\psi}}a_{\mathcal{O}}[n,l+1]C_{l+1}^{(\nu)}(\eta)$
$+\sum_{{\cal\tilde{O}}\in\psi^{\dagger}\times\psi}|x|^{\Delta_{\cal\tilde{O}}-2\Delta_{\psi}-1}\Big{[}(\Delta_{\cal\tilde{O}}-2\Delta_{\psi}-l)a_{{\cal\tilde{O}}_{+}}C^{(\nu)}_{l+1}(\eta)$
$\qquad\qquad\qquad+\Big{(}l+\frac{2l(\Delta_{\cal\tilde{O}}-2\Delta_{\psi}-l)}{(l+1)(l+d-1)}\Big{)}a_{{\cal\tilde{O}}_{-}}C^{(\nu)}_{l-1}(\eta)\Big{]}.$ (33)
Now, again by using (2.1) in the above expression we get
$g_{3}(x)=\sum_{l=1}^{\infty}2|x|^{l-2}a^{f}_{\mathcal{O}}[0,l]C_{l}^{(\nu)}(\eta)+\sum_{n=0}^{\infty}\Big{(}|x|^{2n}(1+\frac{2n}{d})a^{f}_{\mathcal{\tilde{O}}_{-}}[n,0]+|x|^{2n-2}a^{f}_{\cal O}[n,0]\Big{)}C_{0}^{(\nu)}(\eta)+$
$\sum_{\begin{subarray}{c}n=0\\ l=1\end{subarray}}^{\infty}|x|^{2n+l}\left[2a^{f}_{\mathcal{O}}[n+1,l]+2(n+1)a^{f}_{\mathcal{\tilde{O}}_{+}}[n+1,l]+\bigg{(}l+1+\frac{4n(l+1)}{(l+2)(l+d)}\bigg{)}a^{f}_{\mathcal{\tilde{O}}_{-}}[n,l]\right]C_{l}^{(\nu)}(\eta).$ (34)
2.3 Twisted fields
For the theories with chemical potential $\mu$, we should look for the one point functions of conserved currents constructed from scalar $\tilde{\phi}$ and fermion fields $\tilde{\psi}$ with covariant derivative structure as given below
${\cal O}=\tilde{\phi}^{*}D_{\mu_{1}}\cdots D_{\mu_{l}}\tilde{\phi},$
$\tilde{\cal O}_{0}=\tilde{\psi}^{\dagger}D_{\mu_{1}}\cdots D_{\mu_{l}}\tilde{\psi},$
$\tilde{\cal O}_{+}=\tilde{\psi}^{\dagger}\gamma_{\mu_{1}}D_{\mu_{2}}\cdots D_{\mu_{l}}\tilde{\psi},$
$\tilde{\cal O}_{-}=\tilde{\psi}^{\dagger}\gamma^{\mu}D_{\mu}D_{\mu_{1}}\cdots D_{\mu_{l}}\tilde{\psi},$ (35)
where $D_{\mu}=\partial_{\mu}-i\hat{\mu}\delta_{\tau,0}$ and $\hat{\mu}=i\mu$. Note that we will work with imaginary chemical potential $\mu=-i\hat{\mu}$ till we actually find thermal one point functions of the operators given above and finally we analytically continue to real chemical potential of the results.
Thus if one uses the OPE of the scalar field with periodic boundary condition or fermions with anti-periodic boundary condition it is not easy to read out thermal one point functions of such operators with covariant derivatives from the OPE. But with the following redefinition of the fields
$\tilde{\phi}(x)=e^{i\hat{\mu}\tau}\phi(x)\qquad{\rm and}\qquad\tilde{\psi}(x)=e^{i\hat{\mu}\tau}\psi(x),$ (36)
$\tilde{\psi}$ and $\tilde{\phi}$ are called the twisted fields and they satisfy the twisted boundary conditions as given below
$\phi(\tau+1,\vec{x})=e^{-i\hat{\mu}}\phi(\tau,\vec{x})\qquad{\rm and}\qquad\psi(\tau+1,\vec{x})=-e^{-i\hat{\mu}}\psi(\tau,\vec{x}).$ (37)
We can see that the operators in (2.3) can be written in terms of normal derivatives instead of covariant derivatives using the twisted fields.
${\cal O}=\phi^{*}\partial_{\mu_{1}}\cdots\partial_{\mu_{l}}\phi,$
$\tilde{\cal O}_{0}=\psi^{\dagger}\partial_{\mu_{1}}\cdots\partial_{\mu_{l}}\psi,$
$\tilde{\cal O}_{+}=\psi^{\dagger}\gamma^{\mu_{1}}\partial_{\mu_{2}}\cdots D_{\mu_{l}}\psi,$
$\tilde{\cal O}_{-}=\psi^{\dagger}\gamma^{\mu}\partial_{\mu}\partial_{\mu_{1}}\cdots\partial_{\mu_{l}}\psi.$ (38)
Thus we will study the OPEs of the twisted fields where the operators with the covariant derivative structures are manifestly present. Note that from now onward we use $\tilde{\phi}$ and $\tilde{\psi}$ to denote untwisted fields while $\phi$ and $\psi$ denote the twisted fields.
3 Wess-Zumino model with chemical potential($\mu$) and large $\mu$
[…]
$S=\int_{0}^{\beta}d\tau\int d^{2}x\Big{[}(\partial_{\tau}-\mu)\tilde{\phi}_{a}^{*}(\partial_{\tau}+\mu)\tilde{\phi}_{a}+{|\partial_{i}\tilde{\phi}_{a}|^{2}}+\tilde{\psi}_{a}^{\dagger}\gamma_{0}(\partial_{\tau}+\mu)\tilde{\psi}_{a}+{\psi_{a}^{\dagger}\gamma^{i}\partial_{i}\tilde{\psi}_{a}}$
$+\frac{64\lambda^{2}}{N^{2}}(\tilde{\phi}_{a}^{*}\tilde{\phi}_{a})^{3}+\frac{\lambda}{N}\tilde{\psi}_{a}^{\dagger}\tilde{\psi}_{a}\tilde{\phi}_{c}^{*}\tilde{\phi}_{c}+8\frac{\lambda}{N}\tilde{\psi}_{a}^{\dagger}\tilde{\psi}_{b}\tilde{\phi}^{*}_{b}\tilde{\phi}_{a}+\frac{8\lambda}{N}(\tilde{\psi}^{\dagger}_{a}\tilde{\phi}_{a}+\tilde{\psi}_{a}\tilde{\phi}^{*}_{a})^{2}\Big{]},$ (39)
where $a=1,\cdots,N$ and $i=1,2.$ The partition function is given by the following euclidean path integral on $S^{1}_{\beta}\times R^{2}$,
$Z=\int{\cal D}\tilde{\phi}^{*}{\cal D}\tilde{\phi}{\cal D}\tilde{\psi}^{\dagger}{\cal D}\tilde{\psi}e^{-S},$ (40)
The last term involving $(\psi^{\dagger}_{a}\phi_{a}+\psi_{a}\phi_{a})^{2}$ in the action (3) does not contribute to the leading order in large $N$ so we drop this term. Then we apply the standard trick of Hubbard-Stratanovich transformation to simplify the interaction terms by introducing auxiliary fields $\sigma$ and $\zeta$ as shown below. The use of the definition for the Dirac delta as given below,
$1=\int{\cal D}\zeta{\cal D}\sigma e^{i\int\zeta(\sigma-\frac{\phi_{c}^{*}\phi_{c}}{N})},$ (41)
brings the partition function in the following form,
$Z=\int{\cal D}\tilde{\phi}^{*}{\cal D}\tilde{\phi}{\cal D}\tilde{\psi}^{\dagger}{\cal D}\tilde{\psi}{\cal D}\zeta{\cal D}\sigma\exp\Big{[}-\int d^{3}x\Big{(}(\partial_{\tau}-\mu)\tilde{\phi}_{a}^{*}(\partial_{\tau}+\mu)\tilde{\phi}_{a}+{|\partial_{i}\tilde{\phi}_{a}|^{2}}$
$+\tilde{\psi}_{a}^{\dagger}\gamma_{0}(\partial_{\tau}+\mu)\tilde{\psi}_{a}+{\tilde{\psi}_{a}^{\dagger}\gamma^{i}\partial_{i}\tilde{\psi}_{a}}+64\lambda^{2}N\sigma^{3}+8\lambda\sigma\tilde{\psi}_{a}^{\dagger}\tilde{\psi}_{a}-i\zeta\sigma+i\zeta\frac{\tilde{\phi}_{a}^{*}\tilde{\phi}_{a}}{N}\Big{)}\Big{]}.$ (42)
Rescaling $\sigma$ and $\zeta$ in the following manner,
$\sigma=\frac{\sigma^{\prime}}{8\lambda}\qquad{\rm and}\qquad\zeta=-iN\zeta^{\prime}.$ (43)
The partition function becomes,
$Z=\int{\cal D}\tilde{\phi}^{*}{\cal D}\tilde{\phi}{\cal D}\tilde{\psi}^{\dagger}{\cal D}\tilde{\psi}{\cal D}\zeta{\cal D}\sigma\exp\Big{[}-\int d^{3}x\Big{(}(\partial_{\tau}-\mu)\tilde{\phi}_{a}^{*}(\partial_{\tau}+\mu)\tilde{\phi}_{a}+{|\partial_{i}\tilde{\phi}_{a}|^{2}}$
$+\tilde{\psi}_{a}^{\dagger}\gamma_{0}(\partial_{\tau}+\mu)\tilde{\psi}_{a}+{\tilde{\psi}_{a}^{\dagger}\gamma^{i}\partial_{i}\tilde{\psi}_{a}}+\zeta\tilde{\phi}_{a}^{*}\tilde{\phi}_{a}+\sigma\tilde{\psi}_{a}^{\dagger}\tilde{\psi}_{a}+\frac{N}{8\lambda}\sigma^{3}-N\frac{\zeta\sigma}{8\lambda}\Big{)}\Big{]}.$ (44)
Only the zero modes of $\zeta$ and $\sigma$ contribute to the partition function in the leading order in large $N$. Thus ignoring the fluctuations around the zero mode which contribute to subleading order in large $N$, the path integral in the auxiliary fields $\zeta$ and $\sigma$ reduces to ordinary integrals along the variables $\zeta_{0}$ and $\sigma_{0}$ which are zero modes of $\zeta$ and $\sigma$ respectively.
[…]
where
$\log Z_{b}(m_{b},-i\hat{\mu})$ $=-\sum_{n=-\infty}^{\infty}\int\frac{d^{2}p}{(2\pi)^{2}}\log[\big{(}\frac{2\pi n}{\beta}-\hat{\mu}\big{)}^{2}+p^{2}+m_{b}^{2}],$
$=\frac{1}{6\pi\beta^{2}}\Big{(}\beta^{3}m_{b}^{3}+3\beta m_{b}(\text{Li}_{2}(e^{-\beta(m_{b}+i\hat{\mu})})+\text{Li}_{2}(e^{-\beta(m_{b}-i\hat{\mu})}))$
$\qquad+3\text{Li}_{3}(e^{-\beta(m_{b}+i\hat{\mu})})+3\text{Li}_{3}(e^{-\beta(m_{b}-i\hat{\mu})})\Big{)},$ (46)
and,
$\log Z_{f}(m_{f},-i\hat{\mu})$ $=\sum_{n=-\infty}^{\infty}\int\frac{d^{2}p}{(2\pi)^{2}}\log[\big{(}\frac{(2n+1)\pi}{\beta}-\hat{\mu}\big{)}^{2}+p^{2}+m_{f}^{2}],$
$=-\frac{1}{6\pi\beta^{2}}\Big{(}\beta^{3}m_{f}^{3}+3\beta m_{f}(\text{Li}_{2}(-e^{-\beta(m_{f}+i\hat{\mu})})+\text{Li}_{2}(-e^{-\beta(m_{f}-i\hat{\mu})}))$
$\qquad+3\text{Li}_{3}(-e^{-\beta(m_{f}+i\hat{\mu})})+3\text{Li}_{3}(-e^{-\beta(m_{f}-i\hat{\mu})})\Big{)}.$ (47)
The evaluation of (3) and (3) involve regularisation of divergent integrals as discussed in the appendix A. Note that these integrals are evaluated for imaginary chemical potential but here we will use the analytic continuation by taking $\hat{\mu}=i\mu$.
Now the integrals over the zero modes of the auxiliary fields in (45) are evaluated by the saddle point approximation in $\zeta_{0}$ and $\sigma_{0}$. If the saddle point of the integrals are found to be at $\zeta_{0}=m_{b}^{2}$ and $\sigma_{0}=m_{f}$, we have the following pair of saddle point conditions for the integral (45)
$\partial_{\zeta_{0}}\log Z(\zeta_{0},\sigma_{0})|_{\zeta_{0}=m_{b}^{2},\sigma_{0}=m_{f}}=0,\qquad\text{and}\qquad\partial_{\sigma_{0}}\log Z(\zeta_{0},\sigma_{0})|_{\zeta_{0}=m_{b}^{2},\sigma_{0}=m_{f}}=0.$ (48)
These saddle point conditions in $\zeta_{0}$ and $\sigma_{0}$ are called the gap equation and further simplified to give the following equations, with the choice $\beta=1$,
$\frac{m_{f}}{4\lambda}$ $=-\frac{m_{b}}{2\pi}-\frac{1}{2\pi}[\log(1-e^{-m_{b}+\mu})+\log(1-e^{-m_{b}-\mu})],$
$\frac{m_{b}^{2}}{8\lambda}$ $=\frac{3m_{f}^{2}}{8\lambda}+\frac{m_{f}^{2}}{2\pi}+\frac{m_{f}}{2\pi}[\log(1+e^{-m_{f}+\mu})+\log(1+e^{-m_{f}-\mu})].$ (49)
The above pair of coupled transcendental equations are studied numerically at $\mu=0$ in DeWolfe:2019etx . But, at large $\mu$ these equations become analytically tractable and solved in the following manner. Let us first investigate the asymptotic form of the 1st equation from (3) at large $\mu$ as given below, note that as we are looking for real and positive solutions of both $m_{b}$ and $m_{f}$, and we need $10<m_{b}-\mu<$ 1 to have consistent solution for $m_{b}$,
$\frac{\pi m_{f}}{2\lambda}+\log(m_{b}-\mu)+m_{b}=0.$ (50)
Now we can solve for $m_{b}$ from the above equation, and it is given below
$m_{b}$ $=\mu+W\left(e^{-\mu-\frac{\pi m_{f}}{2\lambda}}\right)=\mu+e^{-\mu-\frac{\pi m_{f}}{2\lambda}}+\cdots,$ (51)
where $W(x)$ is product log function. Now we substitute this solution at the leading order in $\mu$ i.e, $m_{b}=\mu$ in the 2nd equation of (3) to obtain
$-\frac{\mu^{2}}{8\lambda}+\frac{3m_{f}^{2}}{8\lambda}+\frac{m_{f}^{2}}{2\pi}+\frac{m_{f}(\mu-m_{f})}{2\pi}=0.$ (52)
Thus solving this we obtain the large $\mu$ asymptotic formula for the thermal mass of fermions to be
$m_{f}=\frac{(\sqrt{4\lambda^{2}+3\pi^{2}}-2\lambda)\mu}{3\pi}+O(e^{-\mu}),$ (53)
keeping only the leading order contribution at large $\mu$. Note that in taking $\mu\to\infty$ limit of (52) we allowed only $m_{f}<\mu$ as there is no consistent solution otherwise.
The above solutions are valid for $\lambda>0.$ From the gap equations it is easy to see that at large $\lambda$, the fermion mass vanishes and the model reduces to a combination of $N$ free fermions and $N$ complex scalars at the non-trivial fixed point or the critical point Filothodoros:2016txa  [^3] In Filothodoros:2016txa , finite imaginary potential was considered which can be analytically continued to real chemical potential. at finite chemical potential with no fermion boson interaction. These large $\mu$ asymptotic solutions are tested against the numerical solutions of the gap equations and shown in the figure 1.
(a) $\lambda=1$
(b) $\lambda=4$
(c) $\lambda=8$
(d) $\lambda$=12
Figure 1: Large $\mu$ asymptotic solutions for the thermal masses given in (51) and (53) are tested against the numerical solution of the gap equations for different values of the coupling constant $\lambda$. The blue and orange curves refer to the numerical solutions for $m_{b}$ and $m_{f}$ of the coupled equations (3). While the green and red straight lines represent the asymptotic formulae for $m_{b}$ and $m_{f}$ respectively at large $\mu$.
4 OPE inversion formula
In this section we apply the OPE inversion formula on both scalar and fermion two point function of the Wess-Zumino model at the non-trivial fixed point of the theory characterised by the gap equations (3). Applying the OPE inversion formula we will evaluate the thermal one point functions of all the different classes of the operators constructed of scalar and fermion bilinears. Note that we will not use the large $\mu$ solutions in this section. We will use the imaginary chemical potential $i\hat{\mu}$ in the analysis of the OPE inversion formula. By writing down the thermal two point of scalars and fermions in the theory at large $N$ in terms of the thermal masses $m_{b}$ and $m_{f}$ satisfying the gap equations (3) and using the OPE inversion formula we will evaluate the thermal one point functions expressed in terms of these thermal masses $m_{b}$ and $m_{f}$. We will study the asymptotic limit for these thermal one point functions at the large $\mu$ in the next section, after analytically continuing these answers for the thermal one point functions to real chemical potential. The study being presented in this section is very similar to the study of inversion formula on the two point functions of the conformal fixed point of $U(N)$ scalars and Gross-Neveu model with chemical potential for the bosonic and fermionic two point functions respectively found in our earlier study for any odd space-time dimension David:2024naf . The only difference here is that we have started with more general structure of the OPEs for the scalar and fermion two point functions given in (2.1) and (2.2). We have allowed the bosonic bilinear operators to arise in the OPE of the fermionic two point functions and vice versa. But as a consequence of the study in this section we will see that the two point function for scalar only admits the scalar bilinears in the OPE at the large $N$ fixed point and so is the fermion two point function allowing only fermionic bilinears in its OPE.
4.1 Scalar two point function
Scalar two point function of the twisted field on $S^{1}\times R^{2}$ at leading order in large $N$ at imaginary chemical potential $\mu=i\hat{\mu}$ is given as the following,
$g^{b}(\tau,\vec{x})=\langle\phi^{*}(\tau,\vec{x})\phi(0,0)\rangle=\sum_{n=-\infty}^{\infty}\int\frac{d^{2}k}{(2\pi)^{2}}\frac{e^{i(\tilde{\omega}_{n}\tau+\vec{k}\cdot\vec{x})}}{\tilde{\omega}_{n}^{2}+\vec{k}^{2}+m_{b}^{2}}.$ (54)
where $\tilde{\omega}_{n}=\frac{2\pi n}{\beta}-\hat{\mu}$. Note that the above correlator is computed from the action (3) using the large $N$ tricks as explained earlier but large $\mu$ limit is not being used. In fact we will not be using large $\mu$ limit in the two point function. Once we obtain the thermal one point function of the higher spin currents by applying the OPE inversion formula on the two point function, we will directly go for large $\mu$ expansions after analytically continuing to real chemical potential by taking $\hat{\mu}=-i\mu$ for thermal one point functions as mentioned earlier.
We recast the above expression as the following with the choice $\beta=1$
$g^{b}(\tau,\vec{x})=\sum_{n=-\infty}^{\infty}e^{-i\tau\hat{\mu}}\int\frac{d\omega d^{2}k}{(2\pi)^{2}}\frac{e^{i(\omega\tau+\vec{k}\cdot\vec{x})}}{(\omega-\hat{\mu})^{2}+\vec{k}^{2}+m_{b}^{2}}\delta(\omega-2\pi n).$ (55)
Now by substituting the Fourier series of the Dirac comb as given below
$\sum_{n\in{\mathbb{Z}}}\delta(\omega-2\pi n)=\frac{1}{2\pi}\sum_{m\in\mathbb{Z}}e^{im\omega},$ (56)
in the (55), we can convert the sum over Matsubara frequencies to the sum over the images in the Euclidean time direction, this method is known as Poisson resummation
$g^{b}(\tau,\vec{x})$ $=\sum_{n=-\infty}^{\infty}e^{in\hat{\mu}}\int\frac{d\omega d^{2}k}{(2\pi)^{3}}\frac{e^{i\omega(\tau+n)+i\vec{k}\cdot\vec{x}}}{\omega^{2}+\vec{k}^{2}+m_{b}^{2}},$
$=\sum_{n=-\infty}^{\infty}e^{in\hat{\mu}}\frac{e^{-m_{b}\sqrt{(\tau+n)^{2}+\vec{x}^{2}}}}{4\pi\sqrt{(\tau+n)^{2}+\vec{x}^{2}}}.$ (57)
Now the two point function of the form (54) has been studied in David:2024naf , to evaluate the thermal one point functions for the model of $U(N)$ scalars at large $N$ with quartic interaction at finite chemical potential using OPE inversion formula at finite temperature. OPE inversion formula is formulated based on the definition of a spectral function $\hat{a}(\Delta,l)$ where $l$ is the spin of the operators which occur in the OPE of the two point function. The spectral function $\hat{a}(\Delta,l)$ admits poles at the dimension of the operators($\Delta_{\mathcal{O}}$) or the operator spectrum in the OPE at a certain value of spin($l$). From the negative residue at these poles at $\Delta=\Delta_{\mathcal{O}}$ the thermal one point functions of spin-$l$ operators of dimension $\Delta_{\mathcal{O}}$ are being read out as shown below,
$a_{\mathcal{O}}[n,l]$ $=-\hat{a}(\Delta,l)|_{{\rm Res\ at\ }\Delta=\Delta_{\mathcal{O}}=2n+l+2\Delta_{\phi}}.$ (58)
Now, we define,
$z=\tau+ix=rw\qquad{\rm and}\qquad\bar{z}=\tau-ix=rw^{-1},$ (59)
where $\tau$ is the imaginary time direction as usual and $x$ is the spatial distance in two dimensional plane. And in the complex $w$-plane the spectral function has contribution due to the branch cut discontinuity of the two point function in the $w$ plane denoted by $\hat{a}_{\rm disc}(\Delta,l)$ and the circular arc at infinity $\hat{a}_{\rm arc}(\Delta,l)$ as given below,
$\hat{a}(\Delta,l)=-\big{(}\hat{a}_{\rm disc}(\Delta,l)+\theta(l_{0}-l)\hat{a}_{\rm arc}(\Delta,l)\big{)},$ (60)
The arc contribution vanishes unless the angular momenta is bounded above, up to a given angular momentum $l_{0}$. Here $\hat{a}_{\rm disc}(\Delta,l)$ is given by,
$\hat{a}_{\rm disc}(\Delta,l)=K_{l}\int_{0}^{1}\frac{d\bar{z}}{\bar{z}}\int_{1}^{1/\bar{z}}\frac{dz}{z}(z\bar{z})^{\Delta_{\phi}-\frac{\Delta}{2}-\nu}(z-\bar{z})^{2\nu}F_{l}\bigg{(}\sqrt{\frac{z}{\bar{z}}}\bigg{)}{\rm Disc}[{g^{b}}(z,\bar{z})],$ (61)
[…]
and,
${\rm Disc}[g^{b}(z,\bar{z})]=\frac{1}{i}\big{(}g^{b}(z+i\epsilon,\bar{z})-g^{b}(z-i\epsilon,\bar{z})\big{)}.$ (62)
While the rest of the contribution $a_{\rm arc}(\Delta,l)$ involves the contribution at infinity in the $z$-plane.
The integration along a circle of infinite radius in the complex $w$ plane is given by
$\hat{a}_{\rm arcs}(\Delta,l)$ $=2K_{l}\int_{0}^{1}\frac{dr}{r^{\Delta+1-2\Delta_{\phi}}}\times\oint\frac{dw}{iw}\lim_{|w|\rightarrow\infty}\left[\Big{(}\frac{w-w^{-1}}{i}\Big{)}^{2\nu}F_{l}(w^{-1})e^{i\pi\nu}{g}^{b}(r,w)\right].$ (63)
Now from the propagator (4.1) and (62) we have
${\rm Disc}[g^{b}(z,\bar{z})]=\frac{1}{2\pi}\sum_{\begin{subarray}{c}m=-\infty\\ m\neq 0\end{subarray}}^{\infty}e^{-im\hat{\mu}}\,\frac{\cos\left(m_{th}\sqrt{(z-m)(m-\bar{z})}\right)}{\sqrt{(z-m)(m-\bar{z})}}.$ (64)
We substitute this expression into the equation (61) to obtain,
$\hat{a}_{\rm disc}(\Delta,l)=\hat{a}_{\rm disc}^{+}(\Delta,l)+\hat{a}_{\rm disc}^{-}(\Delta,l).$ (65)
Where $a^{\pm}_{\rm disc}(\Delta,l)$ is defined so that the contribution due to $m>0$ and $m<0$ keep separated, as given below
$\hat{a}_{\rm disc}^{\pm}(\Delta,l)=\frac{K_{l}}{2\pi}\sum_{m=1}^{\infty}\int_{0}^{1}\frac{d\bar{z}}{\bar{z}}\int_{m}^{{\rm max}(m,\frac{1}{\bar{z}})}\frac{dz}{z}\frac{(z-\bar{z})F_{l}\left(\sqrt{\frac{\bar{z}}{z}}\right)(z\bar{z})^{-\frac{\Delta}{2}}e^{\mp i\hat{\mu}m}}{\sqrt{(z\mp m)(\pm m-\bar{z})}}$
×$\times\cos\left(m_{\rm th}\sqrt{(z\mp m)(\pm m-\bar{z})}\right).$ (66)
It is easy to see that,
$a_{\rm disc}^{-}(\Delta,l)=(-1)^{l}a^{+}_{\rm disc}(\Delta,l)|_{\hat{\mu}\to-\hat{\mu}}.$ (67)
Now using the transformation,
$\bar{z}=z^{\prime}\bar{z}^{\prime}\qquad{\rm and}\qquad z=mz^{\prime},$ (68)
We obtain,
[…]
Now we expand the it at small $\bar{z}$ and keep only the leading order contribution, performing the above integral from $1$ to $\infty$ in $z$ we get
[…]
$=\frac{2K_{l}}{\pi}\sum_{m=1}^{\infty}\frac{e^{-im\hat{\mu}}}{\sqrt{m}}\frac{\sqrt{\pi}2^{\frac{1}{2}-\Delta}m_{b}^{\Delta-\frac{1}{2}}K_{\Delta-\frac{1}{2}}(mm_{b})}{(-\Delta+l+1)\Gamma(\Delta)},$ (70)
${\rm for}\qquad l>0.$
It is clear from the above expression that the spectral function admits pole at $\Delta=l+1$ which is the dimension of the bosonic operators ${\cal O}[0,l]$ and ${\cal\tilde{O}}_{+}[0,l]$ as these two operators were combined in the 2nd term of the OPE (2.1). And the sub-leading order by order corrections in $\bar{z}$ will give the poles arising due to higher twist($n>0$) operators from the combination of the same classes ${\cal O}[n,l]$ and ${\cal\tilde{O}}_{+}[n,l]$ which can be verified straightforwardedly. This pole structure discards the possibility of the presence of the class of ferminic operator ${\cal\tilde{O}}_{0}[n,l]$ in the bosonic OPE for this model at large $N.$ Thus, using this pole structure in (4.1) and the equation (2.1) we get
$a^{b}_{{\cal\tilde{O}}_{0}}[0,l]=0,$ (71)
which implies that
$C_{\phi\phi{\cal\tilde{O}}_{0}[0,l]}=0,\qquad{\rm at\ large\ }N,$ (72)
As the thermal one point functions of the fermionic operator ${\cal\tilde{O}}_{0}[0,l]$ is non-vanishing. We will evaluate thermal one point function of fermionic operators in the next subsection 4.2, by examining the OPEs for the fermion two point functions. The OPE coefficients $C_{\phi\phi{\cal\tilde{O}}_{+}[0,l]}$ and $C_{\phi\phi{\cal\tilde{O}}_{-}[0,l]}$ can be non-zero in principle. But we know that $C_{\phi\phi{\cal\tilde{O}}_{+}[0,l]}=C_{\phi\phi{\cal\tilde{O}}_{-}[0,l]}=0$ at the Gaussian fixed point where $\lambda=0$ or $m_{b}=m_{f}=0$ as well as at the non-trivial fixed point at strong coupling $\lambda\to\infty$. Thus we expect that it is true for all $\lambda$. But we will explicitly show that this is indeed true for $l=1$ and $l=2$ in the section 6 for all $\lambda$. We evaluate the spin-1 and spin-2 current from the partition function of the theory and comparing these against the thermal expectations $a_{\cal O}[0,1]$ and $a_{\cal O}[0,2]$ respectively we can evaluate $C_{\phi\phi{\cal O}[0,1]},C_{\phi\phi{\cal O}[0,2]}$ for arbitrary coupling. We will observe that $C_{\phi\phi{\cal O}[0,1]},C_{\phi\phi{\cal O}[0,2]}$ do not depend on $\lambda$ at the leading order in large $N$ which implies that $C_{\phi\phi{\cal\tilde{O}}_{+}[0,1]}=C_{\phi\phi{\cal\tilde{O}}_{+}[0,2]}=0$ as explained in section 6. We extrapolate this observation for any $l$, i.e., $a^{b}_{{\cal\tilde{O}}_{+}}[0,l]=0$. Finally taking the negative residue of the spectral function obtained at (4.1) and using (65) we evaluate the thermal one point function of the operator ${\cal O}[0,l]$ to be
$a^{b}_{\mathcal{O}}[0,l]$ $=-(a^{+{(0)}}_{\rm disc}+a^{-(0)}_{\rm disc})|_{{\rm Res\ at\ }\Delta=l+1}$
$=\frac{K_{l}}{l!}\sum_{n=0}^{l}\frac{m_{b}^{l-n}(l-n+1)_{2n}}{2^{l+n}n!}\times\Big{(}\text{Li}_{n+1}(e^{-m_{b}-i\hat{\mu}})+(-1)^{l}\text{Li}_{n+1}(e^{-m_{b}+i\hat{\mu}})\Big{)},$ (73)
${\rm for}\ l>0.$
4.2 Fermion two point functions
We consider the two point correlator of the twisted fermion fields as given below,
$\langle\psi_{\alpha}^{\dagger}(x)\psi_{\beta}(0)\rangle=\frac{i}{2}\sum_{k_{0}\in(2n+1)\pi}\int\frac{d^{2}k}{(2\pi)^{2}}\frac{\gamma^{\mu}_{\alpha\beta}\tilde{k}_{\mu}+im_{f}\delta_{\alpha\beta}}{\tilde{k}^{2}+m_{f}^{2}}e^{i\tilde{k}.x}.$ (74)
with $\tilde{k}_{\nu}=k_{\nu}-\hat{\mu}\delta_{\nu 0}$ and $n\in\mathbb{Z}$. $\alpha,\beta$ denoting the spinor indices. Recall that $m_{f}$ is the thermal mass of the fermions satisfying the gap equations (3). Note that in this propagator we keep the chemical potential to be imaginary as given by $\mu=i\hat{\mu}.$
Now we will be studying this fermionic correlator in three different channels as was defined in (2.2) and apply the OPE inversion formula on each of these channels to evaluate the thermal one point functions of three different classes of operators constructed out of fermion bilinears.
$g_{1}(x)$
The expectation value $a_{{\cal O}_{0}}[0,l]$ is contained in the OPE of the correlation function $g_{1}(x)$. Now using (74) and (2.2) we can evaluate $g_{1}{(x)}$ very easily and we get it to be the following
$g_{1}(x)=\langle\psi_{\alpha}^{\dagger}(x)\psi_{\alpha}(0)\rangle=-m_{f}\sum_{\omega_{m}\in(2m+1)\pi}e^{-i\hat{\mu}\tau}\int\frac{d^{2}k}{(2\pi)^{2}}\frac{e^{i\omega_{m}\tau+i\vec{k}\cdot\vec{x}}}{(\omega_{m}-\hat{\mu})^{2}+\vec{k}^{2}+m_{f}^{2}},$ (75)
for $m\in\mathbb{Z}$.
Now to express $g_{1}(x)$ completely in the position space, we recast the above expression as follows
[…]
With the use of the Fourier series of the Dirac comb as
$\sum_{\omega_{m}\in(2m+1)\pi}\delta(\omega-\omega_{m})=\frac{1}{2\pi}\sum_{m\in\mathbb{Z}}e^{im\omega}(-1)^{m},$ (77)
and using the change of variable $\omega^{\prime}=\omega-\hat{\mu}$ in the next step, we obtain,
[…]
Now using the standard method of Poisson resummation we can rewrite the sum over the matsubara frequencies in the above expression as the sum over images as follows
$g_{1}(z,\bar{z})=-m_{f}\sum_{m\in\mathbb{Z}}(-1)^{m}e^{-i\hat{\mu}m}\frac{e^{-m_{f}\sqrt{(m-z)(m-\bar{z})}}}{4\pi\sqrt{(m-z)(m-\bar{z})}},$ (79)
Applying the Euclidean inversion formula (61) on $g_{1}(z,\bar{z})$ given in (79) in the similar manner as shown for the bosonic correlator in the previous subsection we obtain the following the spectral function for the one point function $\hat{a}^{+(0)}(\Delta,l)$.
$\hat{a}^{+(0)}(\Delta,l)=K_{l}m_{f}\sum_{m=1}^{\infty}\frac{(-1)^{m+1}z^{\frac{l+1}{2}}e^{-i\hat{\mu}m}\left(m^{2}z^{3}\right)^{\frac{1}{2}-\frac{\Delta}{2}}\cos\left(mm_{f}\sqrt{z-1}\right)}{\pi\sqrt{z-1}(-\Delta+l+2)}.$ (80)
Taking the negative residue of the above expression for $\hat{a}^{+(0)}(\Delta,l)$ we get
$-\hat{a}^{+(0)}[0,l]|_{\text{Res at }\Delta=l+2}=K_{l}\sum_{m=1}^{\infty}\frac{2^{\frac{1}{2}-l}(-1)^{m+1}m_{f}^{l+\frac{3}{2}}e^{-i\hat{\mu}m}K_{l+\frac{1}{2}}(mm_{f})}{\sqrt{\pi m}\Gamma(l+1)}.$ (81)
Finally using finite sum representation of $K_{l+\frac{1}{2}}(x)$ for $l\in\mathbb{Z}$, we can perform the infinite sum over $m$ to obtain the thermal one point function to be
$a^{f}_{{\cal\tilde{O}}_{0}}[0,l]$ $=-K_{l}\sum_{n=0}^{l}\frac{m_{f}^{l-n+1}(l-n+1)_{2n}\big{(}\text{Li}_{n+1}\left(-e^{-i\mu-m_{f}}\right)+(-1)^{l}\text{Li}_{n+1}\left(-e^{i\mu-m_{f}}\right)\big{)}}{2^{l+n}n!\Gamma(l+1)},$
${\rm for\ }l>0.$ (82)
Note that the pole structure of $\hat{a}^{+(0)}(\Delta,l)$ in (80) clearly rules out the possibility of the presence of the bosonic operator ${\cal O}[0,l]$ in the OPE (27) as $\hat{a}^{+(0)}(\Delta,l)$ has pole only at $\Delta=l+2$ which stands for the fermionic bilinear operators of the kind ${\cal O}_{0}[0,l]$. Thus from this pole structure in (4.2) and the OPE (27) we can conclude that
$a_{\cal O}^{f}[0,l]=0.$ (83)
And now using (26) and using the fact that the thermal one point function of the scalar bilinears denoted by $b_{\cal O}$ is non-zero as we have seen in equation (4.1), we have
$C_{\psi\psi{\cal O}[0,l]}=0,$ (84)
at large $N$.
$g_{2}(x)$
From (74) and (2.2) we can evaluate the correlator $g_{2}(x)$ as the following
$g_{2}(\tau,\vec{x})$ $=\frac{i}{|x|}\sum_{k_{0}\in\pi(2n+1)}\int\frac{d^{2}k}{(2\pi)^{2}}\frac{\tilde{k}_{\mu}x^{\mu}}{\tilde{k}^{2}+m_{th}^{2}}e^{i\tilde{k}x},$
$=\frac{1}{|x|}x^{\mu}\partial_{\mu}\sum_{k_{0}\in\pi(2n+1)}\int\frac{d^{2}k}{(2\pi)^{2}}\frac{e^{i\tilde{k}x}}{\tilde{k}^{2}+m_{th}^{2}}.$ (85)
The method of Poisson re-summation formula as was used for $g_{1}(z,\bar{z})$ transforms the above Matsubara sum into a sum over images to give
$g_{2}(z,\bar{z})=\sum_{m\in\mathbb{Z}}$ $(-1)^{m+1}e^{-im\hat{\mu}}\bigg{[}-\frac{m}{2}\sqrt{\frac{z}{\bar{z}}}-\frac{m}{2}\sqrt{\frac{\bar{z}}{z}}+\sqrt{z\bar{z}}\bigg{]}\times$
$\frac{e^{-m_{f}\sqrt{(m-z)(m-\bar{z})}}(m_{f}\sqrt{(m-z)(m-\bar{z})}+1)}{4\pi[(m-z)(m-\bar{z})]^{3/2}}.$ (86)
Using the method of OPE inversion formula on the correlator $g_{2}(z,\bar{z})$ given above, we can obtain the thermal one point functions of arbitrary spin fermion bilinears. The method has already been implemented in David:2023uya and the result was obtained at the equation (4.26) of David:2023uya for any odd dimension(we put $k=1$ in the result of David:2023uya as the notation used there is $d=2k+1$). Note that the two point function $g_{2}(z,\bar{z})$ has an extra $e^{-im\hat{\mu}}$ compared the correlator studied in equation (4.13) of David:2023uya , thus taking the factor of $e^{-im\hat{\mu}}$ into account to this calculation we have
$a^{f}_{\mathcal{\tilde{O}}_{+}}[n=0,l]$ $=\frac{l}{{\pi\big{(}\frac{1}{2}\big{)}_{l}}}{\,\sum_{n=0}^{l}\frac{m_{f}^{l-n}(l-n+1)_{2n}}{2^{n+l+2}n!}}\left(\text{Li}_{n+1}(-e^{-i\hat{\mu}-m_{f}})+(-1)^{l}\text{Li}_{n+1}(-e^{-m_{f}+i\hat{\mu}})\right),$ (87)
for $l>0.$
We used the fact that $C_{\psi\psi{\cal O}[0,l]}=0$ from (83) i.e., scalar bilinear operators of the kind ${\cal O}[0,l]$ are not present in the OPE of the fermion two point function given in (2).
Applying the inversion formula on the correlator $g_{2}(z,\bar{z})$ we also have obtained the following result in the equation (4.33) of David:2023uya with $d=3$(or $k=1$ as the results obtained there in odd space time dimensions $d=2k+1,\ k\in\mathbb{Z}$),
$a^{f}_{{\cal\tilde{O}}_{+}}[1,l]+\frac{2(l+1)}{(l+2)(l+3)}a^{f}_{{\cal\tilde{O}}_{-}}[0,l]$
$=\frac{(l+2)}{2\pi(3+2l)(\frac{1}{2})_{l}}\left(\frac{m_{f}}{2}\right)^{l+2}\sum_{n=0}^{l}\frac{(l+1-n)_{2n}}{(2m_{f})^{n}n!}({\rm Li}_{n+1}(-e^{-m_{f}-i\hat{\mu}})+(-1)^{l}{\rm Li}_{n+1}(-e^{-m_{f}+i\hat{\mu}})),$
$\qquad\qquad{\rm for}\ \ l>0.$ (88)
Note that here we have an extra $e^{-im\hat{\mu}}$ in (4) compared to the correlator given in the equation (4.13) of David:2023uya . Thus we have accounted for the required modifications due to this $e^{-im\hat{\mu}}$ in the above expression.
$g_{3}(x)$
Using (74) and (2.2) we can write down $g_{3}(x)$ as given below
$g_{3}(x)$ $=\partial^{\mu}\partial_{\mu}\sum_{k_{0}\in(2n+1)\pi}\int\frac{d^{2}k}{(2\pi)^{2}}\frac{e^{i\tilde{k}\cdot x}}{\tilde{k}^{2}+m_{f}^{2}},$
$=\partial^{\mu}\partial_{\mu}\sum_{m\in\mathbb{Z}}(-1)^{m}e^{i\hat{\mu}m}\frac{e^{-m_{f}\sqrt{(m+\tau)^{2}+{\vec{x}}^{2}}}}{4\pi\sqrt{(m+\tau)^{2}+{\vec{x}}^{2}}}.$ (89)
Thus we have
$g_{3}(x)=m_{f}^{2}\sum_{m\in\mathbb{Z}}\frac{(-1)^{m}e^{-m_{f}\sqrt{(m+\tau)^{2}+{\vec{x}}^{2}}+i\hat{\mu}m}}{4\pi\sqrt{(m+\tau)^{2}+{\vec{x}}^{2}}}.$ (90)
Again by using $C_{\psi\psi\cal O}=0$ from (83) the entire analysis of the inversion formula on $g_{3}(x)$ reduces to the computation of the thermal one point functions of fermion bilinears from the $g_{3}(x)$ for the large $N$ Gross-Neveu model performed in David:2023uya , except the fact that one has to keep track of the $e^{im\mu}$ in the correlator (90) compared to the the correlator given in equation (4.33) of David:2023uya with $d=3$(or $k=1$ as the notation used there $d=2k+1$) [^4] Note that there was a typo in the equation (4.33) of David:2023uya , in the denominator of this formula $(k+\frac{1}{2})_{l}$ is not correct, it should be replaced by $(k-\frac{1}{2})_{l}$.. And keeping this $e^{im\mu}$ in that calculation we get
$2a^{f}_{{\cal\tilde{O}}_{+}}[1,l]+(l+1)a^{f}_{{\cal\tilde{O}}_{-}}[0,l]$
$=\frac{1}{\pi(\frac{1}{2})_{l}}\Big{(}\frac{m_{f}}{2}\Big{)}^{l+2}\sum_{n=0}^{l}\frac{(l-n+1)_{2n}}{(2m_{f})^{n}n!}({\rm Li}_{n+1}(-e^{-m_{f}-i\hat{\mu}})+(-1)^{l}{\rm Li}_{n+1}(-e^{-m_{f}+i\hat{\mu}})).$ (91)
${\rm for}\qquad l>0.$
We can solve (4) and (4) to obtain $a_{{\cal O}_{-}}[0,l]$ and $a_{{\cal O}_{-}}[1,l]$, although we do not show the calculation as it is not so illustrative.
Arc Contributions
In David:2023uya ; David:2024naf we have seen that the contribution to the inversion formula due to the circular arc at infinity given in (63) is non-vanishing only for $l=0$ and it vanishes for $l>0$ for all the scalar and fermion correlators. Thus for $l>0$ it is enough to consider disc contribution only.
For $l=0$, it is necessary to evaluate the arc contribution. Here, we demonstrate that the gap equations can be correctly recovered from the thermal one-point functions of the scalar and fermion currents at $l=0$. For the scalar bilinear at $l=0$ or the thermal expectation of $\phi^{2}$ we must include the arc contribution obtained using the formula (63). For scalar operator ${\cal O}[0,0]$ it has been computed in David:2024naf and given by [^5] See equation (2.50) in David:2024naf .
$-\hat{a}_{\rm arcs}(\Delta,0)|_{{\rm Res\ at}\ \Delta=1}=-\frac{m_{b}}{4\pi}.$ (92)
Now combining the arc contribution (92) and the disc contribution given in (4.1) at $l=0$ we have
$a_{\cal O}[0,0]$ $=-(\hat{a}_{\rm disc}+\hat{a}_{\rm arcs})|_{{\rm Res\ at}\ \Delta=l+1},$
$=-\frac{m_{b}}{4\pi}-\frac{1}{4\pi}(\log(1-e^{-i\hat{\mu}-m_{b}})+\log(1-e^{i\hat{\mu}-m_{b}})).$ (93)
Now from the partition function (3) at the saddle point described in (48) we have
$a_{\cal O}[0,0]=\langle\phi^{2}\rangle_{\beta}=\frac{m_{f}}{8\lambda}.$ (94)
Thus using the above equation and (4) we recover the correct gap equation as was given in the 1st line of (3)
$\frac{\pi m_{f}}{2\lambda}+m_{b}+\log(1-e^{-i\hat{\mu}-m_{b}})+\log(1-e^{i\hat{\mu}-m_{b}})=0.$ (95)
For the one point function of the fermion bilinear ${\cal O}_{0}[0,0]$, the arc contribution has also been evaluated in David:2023uya and it is given by  [^6] See equation (4.49) in David:2023uya , note that the presence of the chemical potential $-i\hat{\mu}$ doesn’t alter this answer, we kept track of an overall -ve sign in our correlator $g_{1}(z,\bar{z})$ in (79) compared to that used in David:2023uya .
$-\hat{a}_{\rm arcs}(\Delta,0)|_{{\rm Res\ at}\ \Delta=2}=\frac{m_{f}^{2}}{4\pi}.$ (96)
Combining arc contribution (96) and disc contribution (81) we obtain for
$a_{{\cal O}_{0}}[0,0]$ $=-(\hat{a}_{\rm disc}+\hat{a}_{\rm arcs})|_{{\rm Res\ at}\ \Delta=2},$
$=\frac{m_{f}^{2}}{4\pi}+\frac{m_{f}}{4\pi}\left(\log\left(1+e^{-m_{f}+i\mu}\right)+\log\left(1+e^{-m_{f}-i\mu}\right)\right).$ (97)
From the partition function (3) at the saddle point (48), we also have
$a_{{\cal O}_{0}}[0,0]=2\langle\psi^{\dagger}\psi\rangle_{\beta}=-\frac{3m_{f}^{2}}{8\lambda}+\frac{m_{b}^{2}}{8\lambda}.$ (98)
Note that the factor of $2$ in the above equation arises as we had normalised the 2-point function in (74) with a factor of $\frac{1}{2}$. Substituting (4) in the above equation, we recover the correct gap equation as was given in the 2nd line of (3).
$\frac{m_{f}^{2}}{2\pi}+\frac{m_{f}}{2\pi}(\log\left(1+e^{-m_{f}+i\mu}\right)+\log(1+e^{-m_{f}-i\mu}))+\frac{3m_{f}^{2}}{8\lambda}-\frac{m_{b}^{2}}{8\lambda}=0.$ (99)
This gap equation can similarly be obtained by calculating $a_{{\cal O}_{-}}[0,0]$, though we do not show it here.
5 Thermal one point functions at large $\mu$
In this section, we use the analytic answers for the thermal masses for complex scalars and fermions at large $\mu$ obtained in (51) and (53) respectively in the expressions for the thermal one point functions evaluated in the section 4 using OPE inversion formula and evaluate the systematic expansions for the thermal one point functions at large $\mu$.
Scalar bilinears
(a) $\lambda=1$
(b) $\lambda=4$
(c) $\lambda=8$
(d) $\lambda=12$
Figure 2: We compare the asymptotic formula (5) for $a_{{\cal O}}[0,1]$ against the the numerical value of the $a_{{\cal O}}[0,1]$ obtained from the (5) with the use numerical solutions of the gap equation (3) for different values of the coupling constant $\lambda$. Blue curve denotes the numerical answer while the orange curve stands for the asymptotic formula (5). We kept $\beta=1$ as usual.
(a) $\lambda=1$
(b) $\lambda=4$
(c) $\lambda=8$
(d) $\lambda=12$
Figure 3: We compare the asymptotic formula (5) for $a_{{\cal O}}[0,4]$ against the the numerical value of the $a_{{\cal O}}[0,4]$ obtained from the (5) with the use numerical solutions of the gap equation (3) for different values of the coupling constant $\lambda$. The blue curve denotes the numerical answer while the orange curve stands for the asymptotic formula (5), note that $\beta=1.$
Let us first focus on the thermal one point functions of the bosonic operators ${\cal O}[0,l]$ evaluated at (4.1), analytically continuing back to real chemical potential by using $\hat{\mu}=i\mu$, we have
$a_{\mathcal{O}}[0,l]$ $=\frac{K_{l}}{l!}\sum_{n=0}^{l}\frac{m_{b}^{l-n}(l-n+1)_{2n}}{2^{l+n}n!}\times\Big{(}\text{Li}_{n+1}(e^{-m_{b}+\mu})+(-1)^{l}\text{Li}_{n+1}(e^{-m_{b}-\mu})\Big{)},$
${\rm for}\ l>0.$ (100)
Now, at large $\mu$ we can write down the asymptotic form of the above expression as presented below
$a_{\cal O}[0,l]=\frac{K_{l}}{l!}\Big{[}-\frac{m_{b}^{l}}{2^{l}}\log(m_{b}-\mu)+\sum_{n=1}^{\infty}\frac{m_{b}^{l-n}(l-n+1)_{2n}}{2^{l+n}n!}\zeta(n+1)+O(e^{-\mu})\Big{]}.$ (101)
We replace $\log(m_{b}-\mu)$ in the above expression by using (50) and subsequently with the use of (51) and (53) in the above expression, this admits the following large $\mu$ expansion given as
$\lim_{|\mu\beta|\rightarrow\infty}a_{\cal O}[0,l]=\frac{2^{l-2}\Gamma(l+1)}{\pi\Gamma(2l+1)}\mu^{l+1}\Bigg{[}1+\frac{\sqrt{4\lambda^{2}+3\pi^{2}}-2\lambda}{6\lambda}$
$+\sum_{n=1}^{l}\frac{(l-n+1)_{2n}\zeta(n+1)}{2^{n}n!}\frac{T^{n+1}}{|\mu|^{n+1}}+O(e^{-\beta|\mu|})\Bigg{]}.$ (102)
Numerical tests have been performed to check the agreement of this above asymptotic formula against the value of the thermal one point function computed from the expression (5), e.g. we have presented the comparison between the asymptotic formula and the numerical values in the figure 2 and 3 for $l=1$ and $l=4$ respectively for different values of $\lambda$. by substituting the numerical solution of the gap equations (3). Note that we have reinstated the temperature dependence in the last line. For spin $l=1\ {\rm and}\ l=2$, we have
$a_{\cal O}[0,1]$ $=\frac{\left(\sqrt{4\lambda^{2}+3\pi^{2}}+4\lambda\right)\mu^{2}}{24\pi\lambda}+\frac{\pi T^{2}}{24}+O(e^{-\frac{\mu}{T}}),$ (103)
$a_{\cal O}[0,2]$ $=\frac{\left(\sqrt{4\lambda^{2}+3\pi^{2}}+4\lambda\right)\mu^{3}}{72\pi\lambda}+\frac{1}{24}\pi\mu T^{2}+\frac{T^{3}\zeta(3)}{4\pi}+O(e^{-\frac{\mu}{T}}).$ (104)
Now taking $\lambda\to\infty$ in the above two equations we have
$\lim_{\lambda\to\infty}a_{\cal O}[0,1]$ $=\frac{\mu^{2}}{4\pi}+\frac{\pi T^{2}}{24}+\cdots,$ (105)
$\lim_{\lambda\to\infty}a_{\cal O}[0,2]$ $=\frac{\mu^{3}}{12\pi}+\frac{1}{24}\pi\mu T^{2}+\frac{T^{3}\zeta(3)}{4\pi}+\cdots.$ (106)
Now these expressions agree with the answers for critical $U(N)$ scalars at large $\mu$ given in equations (2.57) of David:2024naf .
Fermion bilinears
Now we can focus on the thermal one point function of the fermion bilinear operator ${\cal O}_{+}[0,l]$ [^7] We are only studying ${\cal O}_{+}[0,l]$ as the thermal expectation of the other two classes of fermion bilinear operators ${\cal O}_{-}[0,l]$ and ${\cal O}_{0}[0,l]$ are directly proportional to the thermal expectation of ${\cal O}_{+}[0,l]$ as shown in David:2023uya . from (87) after analytically continuing to the real chemical potential by $\hat{\mu}=i\mu$.
$a_{\mathcal{\tilde{O}}_{+}}[n=0,l]$ $=\frac{l}{{\pi\big{(}\frac{1}{2}\big{)}_{l}}}{\,\sum_{n=0}^{l}\frac{m_{f}^{l-n}(l-n+1)_{2n}}{2^{n+l+2}n!}}\left(\text{Li}_{n+1}\left(-e^{\mu-m_{f}}\right)+(-1)^{l}\text{Li}_{n+1}\left(-e^{-m_{f}-\mu}\right)\right).$ (107)
Now we study the large $\mu$ behaviour of the above expression using the following identity
${\rm Li}_{s}(e^{2\pi iz})+(-1)^{s}{\rm Li}_{s}(e^{-2\pi iz})=-\frac{(2\pi i)^{s}}{s!}B_{s}(z),$ (108)
[…]
and we obtain
$a_{{\cal\tilde{O}}_{+}}[0,l]=$ $-\frac{lm_{f}^{l}}{2^{l+2}\pi(\frac{1}{2})_{l}}\sum_{n=0}^{l}\frac{(l+1-n)_{2n}}{(2m_{f})^{n}n!}\frac{(2\pi i)^{n+1}}{(n+1)!}\Big{[}B_{n+1}\Big{(}\frac{1}{2}-i\frac{(\mu-m_{f})}{2\pi}\Big{)}+O(e^{-\mu})\Big{]},$ (109)
$B_{n}(x)$ denotes the Bernoulli polynomial. Now by substituting the solution of the gap equation for the fermionic thermal mass $m_{f}$ obtained in (53) with the use of the following notation
$m_{f}=c\mu\qquad{\rm and}\qquad c=\frac{(\sqrt{4\lambda^{2}+3\pi^{2}}-2\lambda)}{3\pi},$ (110)
and using the following identity for the Bernoulli polynomials
$B_{n}(x+y)=\sum_{k=0}^{n}\binom{n}{k}B_{k}(x)y^{n-k},$ (111)
in the equation (109) we have
$a_{{\cal\tilde{O}}_{+}}[0,l]=-\frac{lm_{f}^{l}}{2^{l+2}\pi(\frac{1}{2})_{l}}\sum_{n=0}^{l}\sum_{k=0}^{n+1}(2\pi i)^{k}\frac{(l+1-n)_{2n}}{2^{n}n!(n+1)!}\binom{n+1}{k}B_{k}\Big{(}\frac{1}{2}\Big{)}\frac{(1-c)^{n+1-k}\mu^{1-k}}{c^{n}}.$ (112)
Note that we have neglected the $O(e^{-\mu})$ correction from (109). Now by interchanging the order of the sum over $n$ and $k$ in the above expression we can organise $a_{{\cal\tilde{O}}_{+}}[0,l]$ systematically in orders of $\mu$ at large $\mu$ as given below
$a_{{\cal\tilde{O}}_{+}}[0,l]=-\frac{l(c\mu)^{l}}{\pi 2^{l+2}\left(\frac{1}{2}\right)_{l}}\sum_{k=0}^{l+1}(2\pi i)^{k}B_{k}\left(\frac{1}{2}\right)\mu^{1-k}\sum_{n=k-1}^{l}$ $\frac{\binom{n+1}{k}(1-c)^{n+1-k}(l-n+1)_{2n}}{2^{n}n!(n+1)!c^{n}}.$ (113)
The sum over $n$ can be performed and expressed in terms of the regularised hypergeometric function ${}_{2}\tilde{F}_{1}$ as given below
$a_{{\cal\tilde{O}}_{+}}[0,l]=-\frac{l(c\mu)^{l}}{\pi 2^{l+2}}\sum_{k=0}^{l+1}\frac{(2\pi i)^{k}B_{k}\left(\frac{1}{2}\right)}{\left(\frac{1}{2}\right)_{l}(2c\mu)^{k-1}}\frac{\,{}_{2}\tilde{F}_{1}\left(k-l-1,k+l;k;\frac{c-1}{2c}\right)}{k!(k+l)_{2-2k}}$ (114)
We have tested this asymptotic formula against the values of the thermal one point function (107) obtained using the numerical solution of the gap equations (3), e.g., we show the comparison between the asymptotic answer given above and the numerically obtained value in the figure 4 and 5 for $l=1$ and $l=4$ respectively for different values of $\lambda$. For illustration, from the above formula we explicitly write down the large $\mu$ asymptotic formula for thermal expectation values of the spin-1 and spin-2 current constructed from fermion bilinear as given below after reinstating the temperature dependence
$a_{{\cal\tilde{O}}_{+}}[0,2]=-\frac{\left(3\pi^{2}-4\lambda^{2}+2\lambda\sqrt{4\lambda^{2}+3\pi^{2}}\right)\mu^{3}}{54\pi^{3}}-\frac{\pi\mu T^{2}}{12}+\cdots,$ (115)
and
$a_{{\cal\tilde{O}}_{+}}[0,1]=-\frac{\left(2\lambda\left(\sqrt{4\lambda^{2}+3\pi^{2}}-2\lambda\right)+3\pi^{2}\right)\mu^{2}}{36\pi^{3}}-\frac{\pi T^{2}}{24}+\cdots.$ (116)
We know that at large $\lambda$ these answers should agree with the one point functions of free fermions at large $\mu$ as mentioned earlier, we have verified that it is indeed true. Note that the large $\mu$ asymptotic form of the higher spin currents for both scalars and fermions given in (5) and (113) admits branch cut singularity at $\lambda=\pm\frac{\sqrt{3}}{2}\pi i$ in the analytically continued complex $\lambda$ plane. For imaginary values of $\lambda$ the theory lies on the non-unitary domain.
(a) $\lambda=1$
(b) $\lambda=4$
(c) $\lambda=8$
(d) $\lambda=12$
Figure 4: We compare the asymptotic formula (113) for $a_{{\cal O}_{+}}[0,1]$ against the the numerical value of the $a_{{\cal O}_{+}}[0,1]$ obtained from the (107) with the use numerical solutions of the gap equation (3) for different values of the coupling constant $\lambda$. The blue curve denotes the numerical answer while the orange curve stands for the asymptotic formula (113), note that we kept $\beta=1$.
(a) $\lambda=1$
(b) $\lambda=4$
(c) $\lambda=8$
(d) $\lambda=12$
Figure 5: We compare the asymptotic formula (113) for $a_{{\cal O}_{+}}[0,4]$ against the the numerical value of the $a_{{\cal O}_{+}}[0,4]$ obtained from the (107) with the use numerical solutions of the gap equation (3) for different values of the coupling constant $\lambda$. The blue curve denotes the numerical answer while the orange curve stands for the asymptotic formula (113), note that we kept $\beta=1.$
6 Agreement with the stress tensor and spin-1 current
In this section we discuss the agreement of the stress tensor and spin-1 current evaluated from the partition function with the correct combination of the bosonic and fermionic spin-2 currents and spin-1 currents respectively obtained from the OPE inversion formula. The partition function at large $N$ given in (45) is evaluated at the saddle point is given by
$\log Z$ $=-N\beta V(\frac{m_{f}^{3}}{8\lambda}-\frac{m_{b}^{2}m_{f}}{8\lambda}-\frac{1}{\beta}\log Z_{b}(m_{b},-i\hat{\mu})-\frac{1}{\beta}\log Z_{f}(m_{f},-i\hat{\mu})).$ (117)
After substituting the expression for $\log Z_{b}(m_{b})$ and $\log Z_{f}(m_{f})$ from (3) and (3) respectively in (117), we use the gap equations (3) to obtain $\log Z$ to be
$\log Z=-\frac{NV}{6\pi\beta^{2}}\Big{[}m_{b}^{2}\beta^{2}(\log(1-e^{-m_{b}\beta+i\hat{\mu}\beta})+\log(1-e^{-m_{b}\beta-i\hat{\mu}\beta}))$
$+3m_{b}\beta({\rm Li}_{2}(e^{-m_{b}\beta+i\hat{\mu}\beta})+{\rm Li}_{2}(e^{-\beta m_{b}-i\hat{\mu}\beta}))$
$+{\rm Li}_{3}(e^{-\beta m_{b}\beta+i\hat{\mu}\beta})+{\rm Li}_{3}(e^{-\beta m_{b}\beta-i\hat{\mu}\beta})$
$-m_{f}^{2}\beta^{2}(\log(1+e^{-m_{f}\beta+i\hat{\mu}\beta})+\log(1+e^{-m_{f}\beta-i\hat{\mu}\beta}))$
$-3m_{f}\beta({\rm Li}_{2}(-e^{-m_{f}\beta+i\hat{\mu}\beta})+{\rm Li}_{2}(-e^{-\beta m_{f}-i\hat{\mu}\beta}))$
$-{\rm Li}_{3}(-e^{-\beta m_{f}\beta+i\hat{\mu}\beta})-{\rm Li}_{3}(-e^{-\beta m_{f}\beta-i\hat{\mu}\beta})\Big{]}.$ (118)
Stress tensor
The stress tensor can be evaluated by differentiating the partition function (117) in the following manner
$T_{00}$ $=\partial_{\beta}\log Z|_{\hat{\mu}\beta={\rm constant}},$
$=-\frac{m_{f}^{3}}{8\lambda}+\frac{m_{b}^{2}m_{f}}{8\lambda}+\frac{1}{6\pi\beta^{3}}\Big{[}\beta^{2}m_{b}^{2}(3\log(1-e^{-\beta m_{b}+i\hat{\mu}\beta})+3\log(1-e^{-\beta m_{b}-i\hat{\mu}\beta})+\beta m_{b})$
$-6\beta m_{b}(\text{Li}_{2}(e^{i\hat{\mu}\beta-m_{b}\beta})+\text{Li}_{2}(e^{-m_{b}\beta-i\hat{\mu}\beta}))-6\text{Li}_{3}(e^{i\hat{\mu}\beta-m_{b}\beta})-6\text{Li}_{3}(e^{-m_{b}\beta-i\hat{\mu}\beta})$
$-\beta^{2}m_{f}^{2}(3\log(1+e^{-\beta m_{f}+i\hat{\mu}\beta})+3\log(1+e^{-m_{f}\beta-i\hat{\mu}\beta})+\beta m_{f})$
$+6\beta m_{f}(\text{Li}_{2}(-e^{i\hat{\mu}\beta-m_{f}\beta})+\text{Li}_{2}(-e^{-m_{f}\beta-i\hat{\mu}\beta}))+6\text{Li}_{3}(-e^{i\hat{\mu}\beta-m_{f}\beta})+6\text{Li}_{3}(-e^{-m_{f}\beta-i\hat{\mu}\beta})\Big{]}.$ (119)
Now with the use of the gap equations (3) in (6) we obtain the stress tensor at the conformal fixed point to be
$T_{00}$ $=\frac{1}{3\pi\beta^{3}}\Big{[}\beta^{2}m_{b}^{2}(\log(1-e^{-\beta m_{b}+i\hat{\mu}\beta})+3\log(1-e^{-\beta m_{b}-i\hat{\mu}\beta}))$
$-3\beta m_{b}(\text{Li}_{2}(e^{i\hat{\mu}\beta-m_{b}\beta})+\text{Li}_{2}(e^{-m_{b}\beta-i\hat{\mu}\beta}))-3\text{Li}_{3}(e^{i\hat{\mu}\beta-m_{b}\beta})-3\text{Li}_{3}(e^{-m_{b}\beta-i\hat{\mu}\beta})$
$-\beta^{2}m_{f}^{2}(\log(1+e^{-\beta m_{f}+i\hat{\mu}\beta})+3\log(1+e^{-m_{f}\beta-i\hat{\mu}\beta}))$
$+3\beta m_{f}(\text{Li}_{2}(-e^{i\hat{\mu}\beta-m_{f}\beta})+\text{Li}_{2}(-e^{-m_{f}\beta-i\hat{\mu}\beta}))+3\text{Li}_{3}(-e^{i\hat{\mu}\beta-m_{f}\beta})+3\text{Li}_{3}(-e^{-m_{f}\beta-i\hat{\mu}\beta})\Big{]},$
$\equiv T_{00}^{b}+T_{00}^{f}.$ (120)
$T_{00}^{b}$ and $T_{00}^{f}$ are contributions to the total stress tensor from the scalars and fermions in the action respectively. The terms in (6) involving $m_{b}$ are contained in $T_{00}^{b}$ and the rest is contained in $T_{00}^{f}$.
Comparing (6) and (6) we have, the expected relation between stress tensor and the free energy of a thermal CFT on $S^{1}\times R^{2}$
$T_{00}=-\frac{2\log Z}{N\beta}.$ (121)
Thus the above expression (6) is the stress tensor evaluated from the partition function. Now to obtain the stress tensor of the theory from the thermal one point functions evaluated using OPE inversion we should add the scalar spin-2 current ${\cal O}[0,2]$ from (4.1) and fermion spin-2 current ${\cal O}_{+}[0,2]$ from (87) respectively with the correct normalisations. We use the normalisations such that the bosonic spin-2 currents ${\cal O}[0,2]$ exactly match with the stress tensor for a free complex scalar and fermion spin-2 current ${\cal O}_{+}[0,2]$ matches with the stress tensor for a free fermion when $m_{b}=m_{f}=\mu=0$.
$-4a_{\mathcal{O}}[0,2]+2a_{\mathcal{\tilde{O}}_{+}}[0,2]=\frac{1}{3\pi}\Big{[}m_{b}^{2}(\log(1-e^{i\hat{\mu}-m_{b}})+\log(1-e^{-i\hat{\mu}-m_{b}}))$
$\qquad\qquad-3m_{b}(\text{Li}_{2}(e^{-m_{b}+i\hat{\mu}})+\text{Li}_{2}(e^{-i\hat{\mu}-m_{b}}))-3(\text{Li}_{3}(e^{-m_{b}+i\hat{\mu}})+\text{Li}_{3}(e^{-i\hat{\mu}-m_{b}}))$
$\qquad\qquad-m_{f}^{2}(\log(e^{i\hat{\mu}-m_{f}}+1)+\log(e^{-i\hat{\mu}-m_{f}}+1))$
$+3m_{f}(\text{Li}_{2}(-e^{-m_{f}+i\hat{\mu}})+\text{Li}_{2}(-e^{-i\hat{\mu}-m_{f}}))+3(\text{Li}_{3}(-e^{-m_{f}+i\hat{\mu}})+\text{Li}_{3}(-e^{-i\hat{\mu}-m_{f}}))\Big{]}.$ (122)
To elaborate the scheme of normalisations we use the following answers for the stress tensor of a free massless complex scalar and a fermion respectively at finite temperature in $d=3$
$T_{00}^{\rm free\ boson}=-\frac{2\zeta(3)}{\pi\beta^{3}},\qquad T_{00}^{\rm free\ fermion}=-\frac{3\zeta(3)}{2\pi\beta^{3}}.$ (123)
Now it is easy to see from equation (4.1), by taking the free theory limit $m_{b}=0$ and $\mu=0$, with $\beta=1$.
$-4a_{\cal O}[0,2]|_{m_{b}=0,\mu=0}=T_{00}^{\rm free\ boson}.$ (124)
Similarly by taking $m_{f}=\mu=0$ in the equation (87) we have
$2a_{{\cal\tilde{O}}_{+}}[0,2]|_{m_{f}=0,\mu=0}=T_{00}^{\rm free\ fermion}.$ (125)
Thus the combination $-4a_{\mathcal{O}}[0,2]+2a_{\mathcal{O}_{+}}[0,2]$ agrees with the stress tensor for our model at the free CFT point where the coupling constant $\lambda=0$. We can also see that the expression (6) exactly matches with the stress tensor of the theory (6) at the interacting CFT obtained from the direct calculation of the partition function at $\beta=1$.
$T_{00}=-4a_{\mathcal{O}}[0,2]+2a_{\mathcal{\tilde{O}}_{+}}[0,2].$ (126)
It clearly shows that the OPE coefficients $C_{\phi\phi{\cal O}[0,2]}$ and $C_{\psi\psi{\cal O}_{+}[0,2]}$ at the interacting CFT with any value of $\lambda$ do not alter from that at the free CFT point which was argued during the discussion of the the OPE inversion formula.
Spin-1 Current
We evaluate the spin-1 current in the theory from the partition function (117) by differentiating w.r.t $\hat{\mu}$ as given below
$J=\frac{i}{\beta}\partial_{\hat{\mu}}\log Z,$ (127)
$=\frac{1}{{2\pi\beta^{2}}}\Big{[}\beta m_{b}(\log(1-e^{-\beta m_{b}+i\beta\hat{\mu}})-\log(1-e^{-\beta(m_{b}+i\hat{\mu})}))$
$+\text{Li}_{2}(e^{-\beta(m_{b}+i\hat{\mu})})-\text{Li}_{2}(e^{i\beta\hat{\mu}-m_{b}\beta})+\beta m_{f}(\log(1+e^{-\beta(m_{f}+i\hat{\mu})})-\log(1+e^{-\beta m_{f}+i\beta\hat{\mu}}))$
$-\text{Li}_{2}(-e^{-\beta(m_{f}+i\hat{\mu})})+\text{Li}_{2}(-e^{i\beta\hat{\mu}-m_{f}\beta})\Big{]},$
$\equiv J^{b}+J^{f}.$
The contribution to the total spin-1 current in the theory due to the scalars is given by $J_{b}$ containing the terms involving $m_{b}$. While $J_{f}$ contains the rest of the terms involving $m_{f}$ which give the contribution due to fermions. The spin-1 current from the OPE inversion formula is obtained by adding thermal one point function of scalar spin-1 current ${\cal O}[0,1]$ from (4.1) and fermion spin-1 current ${\cal O}_{+}[0,1]$ from the equation (87)
$a_{\cal O}[0,1]-a_{{\cal\tilde{O}}_{+}}[0,1]=\frac{1}{4\pi}\Big{(}m_{b}\log(1-e^{i\hat{\mu}-m_{b}})-m_{b}\log(1-e^{-i\hat{\mu}-m_{b}})$
$-\text{Li}_{2}(e^{-m_{b}+i\hat{\mu}})+\text{Li}_{2}(e^{-i\hat{\mu}-m_{b}})+m_{f}\log(e^{-i\hat{\mu}-m_{f}}+1)-m_{f}\log(e^{i\hat{\mu}-m_{f}}+1)$
$-\text{Li}_{2}(-e^{-m_{f}-i\hat{\mu}})+\text{Li}_{2}(-e^{i\hat{\mu}-m_{f}})\Big{)},$ (128)
with $\beta=1$ comparing (127) and (6) we have
$a_{\cal O}[0,1]-a_{{\cal\tilde{O}}_{+}}[0,1]=\frac{J}{2}.$ (129)
This is true for all values $m_{b},m_{f}$ and $\mu$, or in other words it is true for the Gaussian fixed point as well as the interacting CFT which shows that the OPE coefficient $C_{\phi\phi{\cal O}[0,1]}$ and $C_{\psi\psi{\cal\tilde{O}}_{+}[0,1]}$ keep fixed across the free CFT as well as the interacting CFT with arbitrary values of coupling.
Vanishing of $C_{\phi\phi{\cal\tilde{O}}_{+}[0,1]},C_{\phi\phi{\cal\tilde{O}}_{+}[0,2]}$
Here we show that the assumption $C_{\phi\phi{\cal\tilde{O}}[0,l]}=0$ used just before (4.1) is indeed true for $l=1$ and $l=2$. Thus we proceed without the assumption that $C_{\phi\phi{\cal\tilde{O}}_{+}[0,l]}=0$, from the equation (4.1) and (2.1) we have
$(a^{b}_{\mathcal{O}}[0,l]+a^{b}_{{\cal\tilde{O}}_{+}}[0,l])$ (130)
$=\frac{K_{l}}{l!}\sum_{n=0}^{l}\frac{m_{b}^{l-n}(l-n+1)_{2n}}{2^{l+n}n!}\times\Big{(}\text{Li}_{n+1}(e^{-m_{b}-i\hat{\mu}})+(-1)^{l}\text{Li}_{n+1}(e^{-m_{b}+i\hat{\mu}})\Big{)},\qquad{\rm for}\ l>0.$
For $l=1$ we have
$(a^{b}_{\mathcal{O}}[0,1]+a^{b}_{{\cal\tilde{O}}_{+}}[0,1])=\frac{1}{4\pi}\Big{(}m_{b}\left(\log(1-e^{i\hat{\mu}-m_{b}})-\log(1-e^{-i\hat{\mu}-m_{b}})\right)$
$-\text{Li}_{2}(e^{-m_{b}+i\hat{\mu}})+\text{Li}_{2}(e^{-i\hat{\mu}-m_{b}})\Big{)},$ (131)
For $l=2$
$(a^{b}_{\mathcal{O}}[0,2]+a^{b}_{{\cal\tilde{O}}_{+}}[0,2])=\frac{1}{{12\pi}}\Big{(}-m_{b}^{2}(\log(1-e^{-i\hat{\mu}-m_{b}})+\log(1-e^{i\hat{\mu}-m_{b}}))$
$+3m_{b}(\text{Li}_{2}(e^{-m_{b}-i\hat{\mu}})+\text{Li}_{2}(e^{i\hat{\mu}-m_{b}}))+3(\text{Li}_{3}(e^{-m_{b}-i\hat{\mu}})+\text{Li}_{3}(e^{i\hat{\mu}-m_{b}}))\Big{)}.$ (132)
Now using the result from the partition function given in (127)
$J_{b}=\frac{1}{2\pi}\Big{(}m_{b}\left(\log(1-e^{i\hat{\mu}-m_{b}})-\log(1-e^{-i\hat{\mu}-m_{b}})\right)-\text{Li}_{2}(e^{-m_{b}+i\hat{\mu}})+\text{Li}_{2}(e^{-i\hat{\mu}-m_{b}})\Big{)},$ (133)
and from (6)
$T_{00}^{b}$ $=-\frac{1}{{3\pi}}\Big{(}-m_{b}^{2}(\log(1-e^{-i\hat{\mu}-m_{b}})+\log(1-e^{i\hat{\mu}-m_{b}}))$
$+3m_{b}(\text{Li}_{2}(e^{-m_{b}-i\hat{\mu}})+\text{Li}_{2}(e^{i\hat{\mu}-m_{b}}))+3(\text{Li}_{3}(e^{-m_{b}-i\hat{\mu}})+\text{Li}_{3}(e^{i\hat{\mu}-m_{b}}))\Big{)},$ (134)
As our starting point was that $a_{\cal O}^{b}[0,l]$ gives the thermal expectation of the spin-$l$ current constructed from the scalar bilinears which for $l=1$ and $2$ is also evaluated from the partition function as $J_{b}$ and $T_{00}^{b}$ as given in above two equations. Thus comparing (6) with (133) and (6) with (6), it is consistent with the fact that
$a_{\cal O}^{b}[0,1]=2J_{b}\qquad{\rm and}\qquad a_{{\cal\tilde{O}}_{+}}^{b}[0,1]=0,$ (135)
$a_{\cal O}^{b}[0,2]=-4T_{00}^{b}\qquad{\rm and}\qquad a_{{\cal\tilde{O}}_{+}}^{b}[0,2]=0.$ (136)
And as $\langle{\cal\tilde{O}}_{+}[0,l]\rangle\neq 0$ we have
$C_{\phi\phi{\cal\tilde{O}}_{+}[0,1]}=C_{\phi\phi{\cal\tilde{O}}_{+}[0,2]}=0.$ (137)
7 Conclusions
We have studied the large $N$ Wess-Zumino model in presence of the chemical potential($\mu$) in $3d$. Considering the Wess-Zumino model for $2N$ real scalars and $2N$ real fermions without chemical potential in the fundamental of $O(2N)$, we construct the $U(N)$ Wess-Zumino model with $N$ complex scalars and $N$ Dirac fermions. We have introduced a chemical potential in this setup at finite temperature. The non-trivial fixed point of the model is obtained at the saddle point of the action in the zero modes of the auxiliary fields used to linearise the action. The saddle point condition is known as the gap equation. We have found the analytically solvable regime of the gap equations in large $N$ Wess-Zumino model in $3d$ at large chemical potential($\mu$) for arbitrary coupling($\lambda$). By using these analytic solutions of the gap equation at large $\mu$, we evaluate the partition function, stress tensor and spin-1 current of the theory as a perturbative expansion at large $\mu$. Applying the OPE inversion formula we also evaluate the higher spin conserved currents of the theory as a perturbative expansion at large $\mu$. As a direct consequence of the study of the OPE inversion formula we find that at the leading order in large $N$ the OPE coefficients $C_{\phi\phi{\cal O}_{0}[0,l]}$ and $C_{\psi\psi{\cal O}[0,l]}$ vanish for all values of $l$ and thus the relevant three point functions also vanish. Although our OPE analysis can not directly conclude whether the coefficients $C_{\phi\phi{\cal O}_{+}[0,l]}$ and $C_{\phi\phi{\cal O}_{-}[0,l]}$ also vanish. But comparing the results obtained from the partition function we have shown that the $C_{\phi\phi{\cal O}_{+}[0,l]}$ indeed vanishes for $l=1,2$. It would be interesting to evaluate these coefficients directly. An interesting problem would be to extend this study on curved backgrounds such as $S^{1}\times S^{2}$, method developed in David:2024pir for $O(N)$ model in this geometry might be useful. The large chemical potential limit may serve as solvable limits in various other related models.
Acknowledgments
I am deeply indebted to Justin David for numerous discussions, which played a crucial role in shaping this work, as well as for his valuable comments on the final draft. I would like to thank Raghu Mahajan, Ashoke Sen and the String theory group at ICTS-TIFR Bengaluru for their insightful comments during my talk there based on my previous works. I would like to thank the organisers and participants of the National String Meeting 2024, IIT Ropar, for providing me the opportunity to present my previous work and attend encouraging sessions on related topics.
Appendix A Partition function in $3d$ for free boson and fermions
In this appendix we show the calculation of the partition function for free bosons and fermions in $d=3$ to obtain the results used in (3) and (3). We will elaborate on the analytic continuation used to obtain these results.
Free complex scalars
We consider the Euclidean action for a complex scalar of mass $m$ at presence of chemical potential $\mu=-i\hat{\mu}$ as given below
$S=\int d^{3}x((D^{\nu}\tilde{\phi})^{*}D_{\nu}\tilde{\phi}+m^{2}\tilde{\phi}^{*}\tilde{\phi}).$ (138)
As defined earlier $D_{\nu}=\partial_{\nu}-i\hat{\mu}\delta_{\nu 0}$. The partition function of this action given by the following Euclidean path integral
$Z_{b}=\int{\cal D}\tilde{\phi}^{*}{\cal D}\tilde{\phi}e^{-S},\qquad{\rm with}$
$\tilde{\phi}(\tau+\beta,\vec{x})=\tilde{\phi}(\tau,\vec{x}),\ \tilde{\phi}^{*}(\tau+\beta,\vec{x})=\tilde{\phi}^{*}(\tau,\vec{x}).$ (139)
The path integral is performed in the standard approach by the use of mode expansion of the free field $\phi(x)$ to give the partition function as
$\log Z_{b}=-\sum_{n=-\infty}^{\infty}\int\frac{d^{2}p}{(2\pi)^{2}}\log\left[\frac{(2\pi n-\hat{\mu}\beta)^{2}}{\beta^{2}}+\vec{p}^{2}+m^{2}\right].$ (140)
Now the sum over Matsubara frequencies can be performed using the following formula for the regulated sum Klebanov:2011uf
$\sum_{n=-\infty}^{\infty}\log\Big{(}\frac{(n+\alpha)^{2}}{q^{2}}+a^{2}\Big{)}=\log[2\cosh(2\pi q|a|)-2\cos(2\pi\alpha)].$ (141)
to obtain
$\log{Z}_{b}$ $=-\frac{1}{\beta^{2}}\int\frac{d^{2}k}{(2\pi)^{2}}\Big{[}\sqrt{\vec{k}^{2}+m_{\rm th}^{2}\beta^{2}}+2{\rm Re}\Big{(}\log\big{[}1-e^{-\sqrt{\vec{k}^{2}+m_{\rm th}^{2}\beta^{2}}+i\beta\hat{\mu}}\big{]}\Big{)}\Big{]}.$ (142)
The 1st term from the above expression is a diverging integral but we regularise it with the use of the analytic continuation of the following standard integral formula as was done in Giombi:2019upv
$-\frac{1}{\beta^{2}}\int\frac{dkk}{(2\pi)}({k}^{2}+m_{\rm th}^{2}\beta^{2})^{-\alpha}\Big{|}_{\alpha=-\frac{1}{2}}=-\frac{\left(\beta^{2}m^{2}\right)^{1-\alpha}}{4\pi(\alpha-1)\beta^{2}}.$ (143)
The integral in the second term from (142) is convergent and can be evaluated using the expansion of the $\log(1+x)$ in small $x$ and resumed back to give the following result
$\log Z_{b}=\frac{\beta m^{3}}{6\pi}+{\rm Re}\frac{\beta m\text{Li}_{2}\left(e^{-\beta(m-i\mu)}\right)+\text{Li}_{3}\left(e^{-\beta(m-i\mu)}\right)}{\pi\beta^{2}}.$ (144)
Free fermions
Considering the Euclidean action for a free massive fermion in $d=3$ with chemical potential $\mu=-i\hat{\mu}$ as given below
$S=\int d^{3}x(\tilde{\psi}^{\dagger}\gamma^{\mu}D_{\mu}\tilde{\psi}-m\tilde{\psi}^{\dagger}\tilde{\psi}).$ (145)
The partition function of this action is given by the following Euclidean path integral
$Z_{b}$ $=\int{\cal D}\tilde{\psi}^{\dagger}{\cal D}\tilde{\psi}e^{-S},\qquad{\rm with}$
$\qquad\tilde{\psi}(\tau+\beta,\vec{x})$ $=-\tilde{\psi}(\tau,\vec{x}),\ \tilde{\psi}^{\dagger}(\tau+\beta,\vec{x})=-\tilde{\psi}^{\dagger}(\tau,\vec{x}).$ (146)
And it is evaluated as
$\log Z_{f}=\sum_{n=-\infty}^{\infty}\int\frac{d^{2}p}{(2\pi)^{2}}\log\left[\frac{((2n+1)\pi-\hat{\mu}\beta)^{2}}{\beta^{2}}+\vec{p}^{2}+m^{2}\right].$ (147)
Performing the Matsubara sum using (141) we get
$\log Z_{f}=\frac{1}{\beta^{2k}}\int\frac{d^{2}p}{(2\pi)^{2}}\left[\sqrt{\vec{p}^{2}+m^{2}\beta}+2{\rm Re}\log\Big{(}1+e^{-\sqrt{\vec{p}^{2}+m^{2}\beta^{2}}+i\hat{\mu}\beta}\big{)}\right].$ (148)
Again using the analytic continuation of the integral (143) we evaluate the above integral to obtain
$\log Z_{f}$ $=-\frac{1}{6\pi\beta^{2}}\Big{(}\beta^{3}m_{f}^{3}+3\beta m_{f}(\text{Li}_{2}(-e^{-\beta(m_{f}+i\hat{\mu})})+\text{Li}_{2}(-e^{-\beta(m_{f}-i\hat{\mu})}))$
$\qquad+3\text{Li}_{3}(-e^{-\beta(m_{f}+i\hat{\mu})})+3\text{Li}_{3}(-e^{-\beta(m_{f}-i\hat{\mu})})\Big{)}.$ (149)